Jefferson Lab HALL A Logbook for September 09 - All EXCEPT Automatic Entries

  • Main INDEX

    Wednesday

  • 292475 09/30/09 22:09 Bo ( RC ) something wrong with R-arm spot++
  • 292474 09/30/09 22:02 adaq Strange spot++ for R-arm run 4664
  • 292471 09/30/09 21:58 deng PVDIS tagger signal check
  • 292470 09/30/09 21:51 adaq spot++ for run 24942
  • 292460 09/30/09 20:36 Bo ( RC ) Experiment Update
  • 292456 09/30/09 19:44 Jack Battery swap on Compton e-det current displays
  • 292446 09/30/09 18:59 Jack (Techs) Right Dipole
  • 292443 09/30/09 18:55 Jack (Techs) Right Arm status
  • 292435 09/30/09 18:24 brads Moller HVs turned OFF
  • 292434 09/30/09 18:15 R. Michaels CRL mod for PVDIS
  • 292430 09/30/09 17:24 ZAFAR AHMED Calibration of cavity monitors.
  • 292429 09/30/09 17:20 brads Plugged copies of Moller target V->F signal into FADC Moller DAQ
  • 292427 09/30/09 16:10 Kai access report (PVDIS purpose)
  • 292426 09/30/09 15:58 R. Michaels online access to PVDIS scalers
  • 292425 09/30/09 15:51 Bo (current Hall status update
  • 292424 09/30/09 15:45 R. Subedi Synch check
  • 292407 09/30/09 14:59 fansler Checklist
  • 292402 09/30/09 14:33 mcanan target_snap_shots_day_shift
  • 292389 09/30/09 12:36 silwal right arm sync messed up, probably scaler acting up
  • 292388 09/30/09 11:59 brads Moller HVs turned on for testing FADC setup
  • 292371 09/30/09 11:16 silwal changed scripts' calls to match the CODA config in the database
  • 292366 09/30/09 10:58 M Friend New Compton FADC Trigger
  • 292364 09/30/09 10:56 M Friend Compton FADC Trigger
  • 292354 09/30/09 10:31 Kai Pan switch of FADC open now
  • 292353 09/30/09 10:29 Kai Pan electronic change (PVDIS purpose)
  • 292336 09/30/09 00:38 cisbani JT valve slightly closed
  • Tuesday

  • 292335 09/29/09 23:53 Navaphon target chart
  • 292334 09/29/09 23:43 navaphon target screen
  • 292319 09/29/09 19:53 E.Chudakov Moller electronics tune
  • 292318 09/29/09 19:45 Jack (Techs) Right Q1
  • 292307 09/29/09 18:45 navaphon target moved
  • 292305 09/29/09 18:42 brads Moller HVs turned OFF
  • 292293 09/29/09 18:08 brads Moller HVs turned on for cosmics test
  • 292292 09/29/09 18:08 R. Subedi access work for pvdis
  • 292291 09/29/09 18:03 brads FADC Moller config update
  • 292284 09/29/09 17:48 dalton HWP and PC voltages for last 2 weeks
  • 292272 09/29/09 16:28 ZAFAR AHMED Flash ADC fro Moller-2
  • 292271 09/29/09 16:21 ZAFAR AHMED Flash ADC for Moller
  • 292261 09/29/09 16:03 Jack (Techs) Right Q3
  • 292258 09/29/09 15:54 deng pvdis left arm signal check results
  • 292230 09/29/09 14:51 mcanan target_screen_shots_day_shift
  • 292228 09/29/09 14:41 M Friend Right HRS/Compton BCM
  • 292220 09/29/09 13:52 paschke run 12778, reduced low beam cut
  • 292204 09/29/09 11:40 A. Camsonne Compton acpa software updated
  • 292186 09/29/09 09:04 D. Parno Tech: N2 level on LHRS Q3 very low
  • 292185 09/29/09 08:54 D. Parno Day shift summary - Sep. 29, 2009
  • 292184 09/29/09 08:24 D. Parno Restricted access
  • 292183 09/29/09 08:24 D. Parno Heidi is ramping down Q2s
  • 292182 09/29/09 08:06 D. Parno Tech: Right Q1 flow red alarm (fixed itself)
  • 292180 09/29/09 07:51 ZAFAR AHMED Half owl shift summary. 4:00 to 8:00
  • 292172 09/29/09 07:06 Yoomin Target to empty
  • 292170 09/29/09 06:38 ZAFAR AHMED Hall A tools.
  • 292166 09/29/09 05:43 ZAFAR AHMED Spot++
  • 292164 09/29/09 05:38 Saha Shift Summary
  • 292161 09/29/09 05:18 ZAFAR AHMED Mark Dalton fixed the Q1.
  • 292159 09/29/09 04:44 ZAFAR AHMED 85uA
  • 292157 09/29/09 04:42 ZAFAR AHMED Q1 is donw.
  • 292156 09/29/09 03:42 Yoomin Run 12778 with R-HRS Q1 Down
  • 292144 09/29/09 03:07 Yoomin spot++
  • 292139 09/29/09 02:57 dwang compton correlation
  • 292135 09/29/09 01:35 deng target tool snap shot 2
  • 292138 09/29/09 01:34 Yoomin BLumi asymN RMS ~1000ppm
  • 292134 09/29/09 01:34 deng target tool snap shot
  • 292131 09/29/09 00:26 jpchen pressure drops due to temperature decreasing
  • 292127 09/29/09 00:01 cisbani Swing Shift Summary
  • 292126 09/29/09 00:01 M Friend Electron only Compton Runs
  • Monday

  • 292125 09/28/09 23:36 M Friend Electron only Compton Runs
  • 292107 09/28/09 21:32 cisbani Halla Tools
  • 292106 09/28/09 21:12 adaq Target screens
  • 292103 09/28/09 20:26 paschke beam burp cut adjusted
  • 292092 09/28/09 19:03 cisbani Raster Size plot
  • 292089 09/28/09 18:26 Apar Half-wave plate position change
  • 292087 09/28/09 18:20 apar Halfwave plate IN, Slug 16
  • 292083 09/28/09 17:40 Kalyan H2 pressure dropping ( still with in limits)
  • 292076 09/28/09 16:01 jpchen Day shift summary
  • 292073 09/28/09 15:44 jpchen BEM screen
  • 292072 09/28/09 15:42 jpchen Compton chart
  • 292071 09/28/09 15:40 jpchen BPM/BCM charts
  • 292070 09/28/09 15:37 jpchen tools screen
  • 292062 09/28/09 15:05 jpchen target screen
  • 292061 09/28/09 15:05 R. Michaels FADC working on ROC2 (I think)
  • 292060 09/28/09 15:04 jpchen target strip charts
  • 292047 09/28/09 13:33 M Friend Compton back to normal
  • 292045 09/28/09 13:17 jpchen spot++
  • 292044 09/28/09 13:13 M Friend Compton HV
  • 292042 09/28/09 13:06 jpchen Compton HV trip
  • 292039 09/28/09 12:17 R. Michaels close JT valve 1 click
  • 292036 09/28/09 11:32 dalton Changed Hall C IA voltage to 6.2
  • 292031 09/28/09 09:52 mcnulty Charge History (85.4C as of 09:30 today). Shift takers please read this.
  • 292030 09/28/09 09:47 jpchen Target: should keep HighPower heater to non-zero
  • 292025 09/28/09 08:08 G.M. Urciuol Owl shift report
  • 292019 09/28/09 06:51 G.M. Urciuol Hall A tools screen
  • 292018 09/28/09 06:50 Yoomin No more "auto_snap" capture?
  • 292013 09/28/09 06:11 Yoomin Target temperature up to 19.36
  • 292004 09/28/09 03:59 Yoomin Target temperature fluctuating
  • 291992 09/28/09 01:00 Yoomin Target Red Alarm
  • 291987 09/28/09 00:16 Yoomin Target screenshot
  • Sunday

  • 291985 09/27/09 23:51 R. Subedi Shift Summary
  • 291978 09/27/09 22:50 R. Subedi ROC31 is dead
  • 291975 09/27/09 22:42 A. Camsonne Compton run 19793 coincidence run test
  • 291972 09/27/09 22:15 Kalyan Target screens
  • 291957 09/27/09 17:57 Kalyan Target IOC rebooted
  • 291956 09/27/09 17:51 dalton Half wave plate out, slug 15
  • 291949 09/27/09 16:28 R. Subedi Left HRS angle setting
  • 291948 09/27/09 16:21 Kalyan Loop 3 temperature flutuating
  • 291946 09/27/09 15:51 xqian Day Shift Summary
  • 291941 09/27/09 14:33 mhuang spot++
  • 291939 09/27/09 14:00 M Friend Compton Laser on Rates Improved
  • 291933 09/27/09 12:44 M Friend Low Compton Laser on Rates
  • 291930 09/27/09 11:56 mcnulty Charge History (79.9C at 09:30 today)
  • 291929 09/27/09 11:38 mcnulty Harp scan plots from IHA1H03B
  • 291928 09/27/09 11:36 mcnulty Harp scan plots IHA1H03A
  • 291927 09/27/09 11:23 Apar lumi width fine, beam spot ~ 70x70um from the harp scan
  • 291925 09/27/09 11:12 mcnulty Larm s0 HV's off
  • 291924 09/27/09 11:10 mcnulty spot++ on Hydrogen
  • 291921 09/27/09 11:06 mcnulty harp scans done
  • 291920 09/27/09 10:57 mcnulty spot++ (a closer look)
  • 291919 09/27/09 10:52 mcnulty spot++ on Carbon hole 10uA rastered
  • 291914 09/27/09 10:47 mhuang target movement
  • 291913 09/27/09 10:47 mcnulty Target move error
  • 291911 09/27/09 10:35 mcnulty Moving to Carbon hole target
  • 291910 09/27/09 10:32 mcnulty spot++ after -10% change
  • 291908 09/27/09 10:28 silwal dithering plots @ -10 % y dispersion change
  • 291905 09/27/09 10:15 mcnulty Larm s0 HVs on
  • 291904 09/27/09 10:14 silwal dithering plots before changes
  • 291903 09/27/09 10:13 mcnulty Setting up for -10% (moller quad change) phase advance
  • 291898 09/27/09 09:29 mhuang target charts
  • 291895 09/27/09 08:05 mhuang spot++
  • 291889 09/27/09 07:50 ZAFAR AHMED Owl shift summary
  • 291886 09/27/09 07:02 ZAFAR AHMED Beam is down for next 15/20 min.
  • 291884 09/27/09 06:29 ZAFAR AHMED spot++ is normal now
  • 291882 09/27/09 06:26 ZAFAR AHMED spot++ (unusual)
  • 291879 09/27/09 06:10 ZAFAR AHMED ROC 21 crashed.
  • 291878 09/27/09 06:09 dwang correlation plane B
  • 291869 09/27/09 04:08 ZAFAR AHMED spot++
  • 291862 09/27/09 02:46 G.M. Urciuol Target screen snap shot
  • 291861 09/27/09 02:41 G.M. Urciuol Target strip charts
  • 291860 09/27/09 02:28 ZAFAR AHMED Current chart.
  • 291859 09/27/09 02:27 ZAFAR AHMED Hall A tools
  • 291857 09/27/09 02:11 ZAFAR AHMED spot++
  • 291853 09/27/09 01:15 A. Camsonne Test coincidence window for Compton
  • 291850 09/27/09 00:52 AHMED HALLA:p and HALLA:dpp are working now.
  • 291849 09/27/09 00:29 ahmed HALLA:p and HALLA:dpp are not available
  • 291847 09/27/09 00:14 ZAFAR AHMED spot++
  • 291842 09/27/09 00:01 R. Subedi Shift summary
  • Saturday

  • 291837 09/26/09 22:50 R. Subedi dp/p strip chart flat
  • 291836 09/26/09 22:04 Abdurahim Target strip charts
  • 291835 09/26/09 22:03 Abdurahim Target screen snap shot
  • 0 09/26/09 20:58 no_name no_key
  • 291819 09/26/09 19:26 dalton Reboot QWeak crate
  • 291817 09/26/09 18:45 dalton Hall C charge asymmetry
  • 291814 09/26/09 18:29 Abdurahim Target strip charts
  • 291813 09/26/09 18:26 Abdurahim Target screen snapshot
  • 291797 09/26/09 16:33 Abdurahim Target iocha13 rebooted
  • 291793 09/26/09 16:00 A. Deur Day shift summary
  • 291792 09/26/09 15:20 Meddi junk run 12705
  • 291786 09/26/09 14:44 mcnulty Charge History (74.4C at 09:30 today)
  • 291785 09/26/09 14:21 mcnulty Production running resumed
  • 291784 09/26/09 14:18 silwal spot++
  • 291780 09/26/09 14:09 mcnulty beam conditions returned nicely after openinf slits (as expected)
  • 291779 09/26/09 14:04 silwal moller quad @ 20 % change
  • 291778 09/26/09 14:03 silwal moller quad @ 20 % change
  • 291777 09/26/09 14:01 silwal dithering plots @ 10 % moller quad change
  • 291776 09/26/09 14:01 mcnulty Larm s0's HV off
  • 291775 09/26/09 13:59 silwal dithering plots before Moller quad changes
  • 291774 09/26/09 13:58 mcnulty spot++ after changing quad by -10%
  • 291770 09/26/09 13:40 mcnulty Preparing for dithering phase advance (part3)
  • 291767 09/26/09 13:24 mcnulty spot++ after low current running/MCC ioc trip
  • 291764 09/26/09 13:22 A. Camsonne Resuming 100 uA
  • 291753 09/26/09 12:13 A. Camsonne compton rates with slit closed to 10 uA
  • 291752 09/26/09 11:59 mhuang target red alarm
  • 291750 09/26/09 11:52 A. Camsonne Requesting current to 10 uA for Compton test with closing the slit
  • 291749 09/26/09 11:47 A. Camsonne Requesting current to 10 uA for Compton test
  • 291743 09/26/09 11:22 mhuang target red alarm
  • 291742 09/26/09 11:00 A. Camsonne Test ETROC coincidence output sent to FADC
  • 291741 09/26/09 10:34 A. Deur cryotarget Alarm
  • 291736 09/26/09 10:07 mhuang target red alarm
  • 291726 09/26/09 08:54 A. Deur snapshot of BEM
  • 291725 09/26/09 08:50 A. Deur BPM strip tool
  • 291724 09/26/09 08:49 xqian Target screen capture
  • 291722 09/26/09 08:40 A. Deur Hall A tools
  • 291718 09/26/09 08:19 A. Deur raster size check
  • 291714 09/26/09 07:57 xqian Owl shift summary
  • 291711 09/26/09 07:14 xyan target screen snapshots
  • 291710 09/26/09 07:10 Yoomin Sorry, it's IN
  • 291709 09/26/09 07:10 Yoomin Sorry, it's IN
  • 291707 09/26/09 07:05 xyan spot++
  • 291705 09/26/09 06:39 Yoomin IHWP OUT run 12694
  • 291702 09/26/09 06:38 apar Halfwave plate IN, Slug 14
  • 291704 09/26/09 06:37 Yoomin IHWP flipped OUT->IN
  • 291698 09/26/09 04:57 Yoomin Compton DAQ crashed
  • 291690 09/26/09 04:40 xinhu yan spot++
  • 291685 09/26/09 03:12 Yoomin Compton ROC1, ROC2
  • 291684 09/26/09 02:59 Yoomin spot++
  • 291681 09/26/09 02:42 Yoomin CODA reboot
  • 291677 09/26/09 02:29 Xinhu Yan Open JT Valve
  • 291671 09/26/09 01:23 Xinhu Yan Spot++
  • 291662 09/26/09 00:19 Xinhu Yan Target Charts Check
  • 291660 09/26/09 00:00 G.M. Urciuol Swing shift report
  • Friday

  • 291655 09/25/09 23:20 mhuang target strip charts
  • 291654 09/25/09 23:18 mhuang target snapshot
  • 291624 09/25/09 17:47 dalton Prompt output
  • 291613 09/25/09 16:02 Saha Shift Summary
  • 291610 09/25/09 14:47 mcnulty all off in Larm HV slot2 except ch9 and 11
  • 291609 09/25/09 14:46 LeRose Target Screens
  • 291608 09/25/09 14:46 mcnulty Lumi widths just dropped to 1200ppm
  • 291607 09/25/09 14:42 mcnulty Lumi widths are 1400ppm!
  • 291605 09/25/09 14:35 silwal dithering plots, @ 30 % change in last moller quad
  • 291603 09/25/09 14:35 mcnulty Larm s0 HV's off (resuming production)
  • 291602 09/25/09 14:34 LeRose Target Moves
  • 291601 09/25/09 14:34 mcnulty spot++ after phase-advance backout
  • 291599 09/25/09 14:28 mcnulty the rates in the Larm s0 are ok.
  • 291598 09/25/09 14:24 mcnulty spot++ after 30% change in quad
  • 291595 09/25/09 14:17 mcnulty spot++ after initial rty with phase advance (part2)
  • 291594 09/25/09 14:16 Saha BCM/BPM Striptool
  • 291593 09/25/09 14:15 Saha Hall A Tools
  • 291592 09/25/09 14:15 mcnulty SOMEBODY TURNED ON Larm s0 HV
  • 291585 09/25/09 13:45 mcnulty dithering phase-advance setup (part2)
  • 291576 09/25/09 12:16 A. Camsonne Background electron detector
  • 291573 09/25/09 11:05 Kowalski DAQ
  • 291564 09/25/09 10:00 kowalski Spot++
  • 291553 09/25/09 08:26 LeRose Target Screens
  • 291551 09/25/09 07:59 R. Michaels owl shift summary
  • 291548 09/25/09 07:05 R. Michaels getruninfo updated
  • 291547 09/25/09 06:52 R. Michaels updated runstart.tcl for HRS
  • 291545 09/25/09 06:51 reimer target screens
  • 291538 09/25/09 05:00 R. Michaels IHWP out - slug 13
  • 291532 09/25/09 04:54 Yoomin spot++
  • 291529 09/25/09 04:42 R. Michaels switched to disk4 for data
  • 291528 09/25/09 04:40 Yoomin IHWP OUT
  • 291526 09/25/09 04:34 apar Halfwave plate OUT, Slug 13
  • 291524 09/25/09 04:27 reimer Target Temperature
  • 291521 09/25/09 02:49 R. Michaels end CompSimple run 19689
  • 291518 09/25/09 02:04 R. Michaels data checks: run 19689
  • 291517 09/25/09 01:48 Yoomin Lumi RMS
  • 291516 09/25/09 01:47 R. Michaels CompSimple run 19689
  • 291515 09/25/09 01:45 Yoomin diff_bpm10x not found
  • 291514 09/25/09 01:44 Yoomin spot++
  • 291509 09/25/09 01:04 R. Michaels runbird false alarm ?
  • 291508 09/25/09 00:55 R. Michaels nuke parentlocks for firefox
  • Thursday

  • 291499 09/24/09 23:56 jpchen swing shift summary
  • 291497 09/24/09 23:45 jpchen BCM/BPM.
  • 291493 09/24/09 23:34 silwal PITA scan with IHWP IN
  • 291488 09/24/09 22:49 M Friend Changed Right HRS/Compton BCM
  • 291481 09/24/09 21:25 mhuang spot++
  • 291472 09/24/09 20:04 jpchen Compton
  • 291471 09/24/09 20:02 jpchen BCM/BPM
  • 291470 09/24/09 20:01 jpchen tools screen
  • 291467 09/24/09 19:29 Navaphon target charts
  • 291466 09/24/09 19:27 navaphon target screen
  • 291464 09/24/09 19:24 mhuang spot++
  • 291448 09/24/09 15:59 Kowalski Shift Summary
  • 291447 09/24/09 15:42 reimer target screens
  • 291446 09/24/09 15:40 kowalski happex
  • 291444 09/24/09 15:14 mcnulty spot++ on Hydrogen target
  • 291438 09/24/09 15:11 reimer target LH2
  • 291437 09/24/09 15:09 mcnulty spot++ again after y steering
  • 291435 09/24/09 15:05 mercado BCM noise lowered dramatically in slug 11
  • 291434 09/24/09 14:55 mcnulty Charge Time History plot
  • 291431 09/24/09 14:48 mcnulty spot++ on holey Carbon
  • 291428 09/24/09 14:41 mcnulty emerging from phase advance (part1) backout
  • 291427 09/24/09 14:40 reimer Target C hole
  • 291426 09/24/09 14:29 M Friend Compton rates
  • 291425 09/24/09 14:04 reimer target optics
  • 291424 09/24/09 13:54 silwal boiling width ~ 2400 at new quad settings, phase advance looks good
  • 291422 09/24/09 13:30 mcnulty spot++ on Hydrogen (raster current scan)
  • 291418 09/24/09 13:26 mcnulty spot++ after steering in x (part 2)
  • 291415 09/24/09 13:12 gjin Left momentum changed into 1.221 GeV since run 1463
  • 291413 09/24/09 13:10 mcnulty spot++ after steering in x (part 1)
  • 291410 09/24/09 13:04 mcnulty spot++ after MCC steering
  • 291402 09/24/09 12:49 mcnulty spot++ on Holey Carbon
  • 291397 09/24/09 12:39 mcnulty spot++ after phase advance
  • 291395 09/24/09 12:36 reimer target carbon hole
  • 291385 09/24/09 11:26 reimer Target on Empty
  • 291381 09/24/09 11:02 dalton Epics values corrupted during day and swing shift
  • 291380 09/24/09 10:23 dalton Compton restarted
  • 291371 09/24/09 08:45 silwal epics variable readout difficulties
  • 291367 09/24/09 07:58 X. Qian Owl shift summary
  • 291357 09/24/09 05:32 Yoomin spot++
  • 291353 09/24/09 05:14 D. Parno Target tools -- owl shift
  • 291352 09/24/09 05:07 A. Camsonne Bleedthrough / correlation hall C current and Compton central crystal
  • 291324 09/24/09 00:44 A. Camsonne Compton Electron detector to garage position
  • 291321 09/24/09 00:36 A. Camsonne Resumed production
  • 291313 09/24/09 00:02 G.M. Urciuol swing shift report
  • Wednesday

  • 291311 09/23/09 23:57 A. Camsonne Compton software coincidence set to 30 with offset of 20
  • 291310 09/23/09 23:44 A. Camsonne Compton rates are good
  • 291308 09/23/09 23:30 R. Michaels run 12617 and 12618 at 10 uA
  • 291306 09/23/09 23:08 R. Michaels spot++ run 24842
  • 291304 09/23/09 23:03 mcnulty Returning to production (HAPPEx DAQ) mode
  • 291303 09/23/09 22:59 R. Michaels 10 uA Qsq run.
  • 291294 09/23/09 22:44 R. Michaels more runs (waiting for Compton)
  • 291285 09/23/09 22:19 Bob / Dustin more 5 uA running
  • 291276 09/23/09 21:59 A. Camsonne Compton electron detector in position 2mm
  • 291269 09/23/09 21:37 Bob / Dustin Qsq runs at 5 uA.
  • 291254 09/23/09 21:20 Bob/Dustin Qsq at 1 uA. Runs 24830 / 4613
  • 291246 09/23/09 21:03 Bob / Dustin spectrometer DAQ runs
  • 291242 09/23/09 20:58 mcnulty spot++
  • 291239 09/23/09 20:54 A. Camsonne Compton rates with Compton lock
  • 291228 09/23/09 20:49 mcnulty Moving to Holey Carbon Target, do spot++
  • 291227 09/23/09 20:47 A. Camsonne Compton rates after Moller back out
  • 291226 09/23/09 20:45 mcnulty MCC performing Harp scan
  • 291219 09/23/09 20:18 mcnulty Setting up for Qsq measurement
  • 291218 09/23/09 20:14 mcnulty MCC restoring beam after Meoller
  • 291217 09/23/09 20:11 A. Camsonne GSO crystal pulser trigger replaced by photon electron detector trigger output
  • 291216 09/23/09 20:04 O.Glamazdin Moller results
  • 291215 09/23/09 19:51 A. Camsonne Compton electron detector out for Moller back out tune
  • 291214 09/23/09 18:17 silwal script added to change Injector.datamap when switching between Happex and HapNoQweak CODA config, no need to manually modify Injector.datamap
  • 291213 09/23/09 18:15 A. Camsonne Compton electron detector in
  • 291204 09/23/09 17:15 A. Giusa grab of target screens.
  • 291190 09/23/09 16:35 A. Giusa target moving error
  • 291184 09/23/09 16:28 A. Camsonne Compton electron detector out for Moller tuning
  • 291178 09/23/09 16:00 D. Parno Day Shift Summary (23 Sep 2009)
  • 291175 09/23/09 14:27 D. Parno Beamline HV crate tripped again
  • 291173 09/23/09 14:19 mercado CODA restarted
  • 291172 09/23/09 14:07 D. Parno Hall A Tools screen capture
  • 291171 09/23/09 14:06 reimer Target Screens
  • 291170 09/23/09 13:59 D. Parno LHRS N2 levels
  • 291169 09/23/09 13:58 D. Parno dpp fluctuations
  • 291168 09/23/09 13:54 reimer target heater alarm
  • 291165 09/23/09 13:42 dalton Compton cavity optimised
  • 291164 09/23/09 13:27 dalton Compton epics reboot
  • 291156 09/23/09 12:30 D. Parno spot++
  • 291153 09/23/09 12:28 mercado Added bcm7 to Injector datamap
  • 291152 09/23/09 12:21 D. Parno Beamline HV crate tripped
  • 291151 09/23/09 12:16 silwal modified .db files to reflect datamap changes in bpm8/14, and scanclean/scandata flags
  • 291150 09/23/09 12:15 A. Camsonne Added FADC flag in Compton CODA file
  • 291149 09/23/09 12:09 D. Parno Compton ion chamber trips
  • 291146 09/23/09 11:18 dalton Compton DAQ restart
  • 291145 09/23/09 10:56 D. Parno dpp fluctuations ok?
  • 291143 09/23/09 10:40 D. Parno Tiefenback dp/p fluctuations
  • 291142 09/23/09 10:33 mcnulty Charge accumulated so far: 61.8 Coulombs
  • 291139 09/23/09 09:44 D. Parno Beam is back
  • 291136 09/23/09 09:14 D. Parno Beam down
  • 291133 09/23/09 08:55 kowalski DAQ problems
  • 291127 09/23/09 08:32 D. Parno LHRS Momentum Is Correct Again
  • 291124 09/23/09 08:29 D. Parno LHRS Magnet Cycling Instructions Incorrect
  • 291123 09/23/09 08:09 D. Parno Lost communication with iocha14
  • 291122 09/23/09 08:04 ZAFAR AHMED Owl shift summary
  • 291120 09/23/09 07:45 A. Camsonne Rebooting Compton EPICS
  • 291116 09/23/09 06:18 Ahmed Hall A Tools
  • 291114 09/23/09 06:00 AHMED spot++
  • 291109 09/23/09 04:50 AHMED BPM 12 X
  • 291108 09/23/09 04:35 ZAFAR AHMED Compton DAQ
  • 291099 09/23/09 04:13 ZAFAR AHMED BPM 12X
  • 291098 09/23/09 04:08 ZAFAR AHMED BPM12X-Called RC
  • 291097 09/23/09 03:53 ZAFAR AHMED spot++
  • 291093 09/23/09 03:32 ZAFAR AHMED BPM 12X - more than 100K
  • 291090 09/23/09 02:29 deng snap shot of target screen
  • 291089 09/23/09 02:28 deng snap shot of target tools
  • 291086 09/23/09 01:58 Bellini spot++
  • 291081 09/23/09 01:28 ZAFAR AHMED Half Wave Plate = IN
  • 291079 09/23/09 01:25 apar Halfwave plate IN, Slug 12
  • 291077 09/23/09 00:43 ZAFAR AHMED Swing shift summary(09-22-2009)
  • 291076 09/23/09 00:36 AHMED spot++
  • Tuesday

  • 291064 09/22/09 23:25 dwang target screen
  • 291053 09/22/09 20:50 paschke mapping bpms: we have 8 and 14 from iocse9
  • 291052 09/22/09 20:41 paschke, sil Left HRS LED testing
  • 291015 09/22/09 16:00 Saha Shift Summary
  • 291008 09/22/09 13:03 Saha BCM/BPM Striptool
  • 291007 09/22/09 13:01 Saha Hall A Tools Run #12572
  • 291006 09/22/09 13:01 reimer Target screens
  • 291005 09/22/09 12:56 reimer spot++
  • 290991 09/22/09 10:00 A. Camsonne Compton Electron detector current
  • 290980 09/22/09 08:10 G.M Urciuoli owl shit report
  • 290978 09/22/09 07:21 G.M. Urciuol Compton DAQ problems
  • 290977 09/22/09 06:46 A. Giusa Grab of target screens
  • 290972 09/22/09 05:05 Meddi DAQ preblem
  • 290960 09/22/09 03:44 Meddi Compton DAQ problem
  • 290959 09/22/09 03:28 A. Giusa Grab of target screens
  • 290950 09/22/09 01:12 silwal LHRS, can't telnet into ROC3, CODA runs fine, spot++ looks normal
  • 290940 09/22/09 00:14 jpchen alarm Hi limit on ESR 15K supply raised to 25.5 g/s
  • 290939 09/22/09 00:11 jpchen raised alarm limit for fan to 76 Hz
  • 290938 09/22/09 00:02 D. Parno Swing Shift Summary -- September 21, 2009
  • Monday

  • 290934 09/21/09 23:47 bellini Reducing 1% JT valve
  • 290933 09/21/09 23:35 bellini High level alarm changed without reporting
  • 290932 09/21/09 23:09 bellini Yellow fan alarm on L3
  • 290931 09/21/09 23:05 D. Parno Intermittent R_MAG alarm
  • 290928 09/21/09 22:54 bellini Target snapshot at the end of the shift
  • 290927 09/21/09 22:46 D. Parno Hall A tools - snapshot
  • 290926 09/21/09 22:44 D. Parno Changed Compton config to photons only
  • 290925 09/21/09 22:32 D. Parno spot++
  • 290922 09/21/09 22:28 Bellini Another Yellow alarm for fan frequency
  • 290921 09/21/09 22:05 D. Parno 40 minutes without a beam trip
  • 290911 09/21/09 20:17 Bellini Yellow alarm for fan frequency
  • 290908 09/21/09 19:38 D. Parno spot++
  • 290905 09/21/09 19:14 D. Parno IHWP->OUT
  • 290903 09/21/09 19:08 apar Halfwave plate OUT, Slug 11
  • 290901 09/21/09 18:54 Bellini Opening 0.5% more JT valve on LT3
  • 290898 09/21/09 17:58 Bellini Opening 0.5% more JT valve on LT3
  • 290897 09/21/09 17:27 Bellini Opening 1% more JT valve on L3
  • 290892 09/21/09 16:34 bellini Opening 1% JT valve on L3
  • 290884 09/21/09 16:04 mercado Day Shift Summary
  • 290876 09/21/09 15:01 paschke beam trip problem solved, raster at 10x10 setting
  • 290875 09/21/09 14:49 mercado JT Valve adjusted again.
  • 290873 09/21/09 14:29 mercado JT Valve adjustment
  • 290867 09/21/09 13:53 M Friend Compton FADC DAQ config change
  • 290866 09/21/09 13:07 mcnulty Run Coordinator Daily Summary (2009Sep20 - 2009Sep21)
  • 290865 09/21/09 13:06 mercado Raster Changed
  • 290855 09/21/09 12:24 Kai target alarms again
  • 290852 09/21/09 11:31 M Friend Compton network problem
  • 290851 09/21/09 11:01 mcnulty Accumulated charge: 74.06 C
  • 290850 09/21/09 10:40 mcnulty Dithering on but not energizing coils.
  • 290847 09/21/09 10:36 Kai spot++ study
  • 290844 09/21/09 10:21 mcnulty Plan to understand the excesive ep ion chamber trips during the night
  • 290841 09/21/09 09:18 mercado Dithering ON
  • 290840 09/21/09 09:04 mcnulty Dithering turned back on per request of MCC
  • 290839 09/21/09 08:48 mercado Tuning Beam
  • 290831 09/21/09 08:11 Yoomin Owl Shift Summary
  • 290827 09/21/09 07:44 mcnulty Turned Dithering off
  • 290826 09/21/09 07:43 Yoomin Beam status during this shift 2
  • 290825 09/21/09 07:40 Yoomin Beam status during this shift 1
  • 290818 09/21/09 06:32 Yoomin spot++
  • 290817 09/21/09 06:31 dwang target status
  • 290812 09/21/09 05:57 Yoomin Configuration changed to HapNoQweak
  • 290806 09/21/09 05:29 Yoomin About the unstable beam
  • 290803 09/21/09 04:54 Yoomin Unstable beam
  • 290802 09/21/09 04:53 dwang target strip charts
  • 290793 09/21/09 03:51 Yoomin Unstable beam
  • 290787 09/21/09 02:25 Yoomin Compton HV trip
  • 290784 09/21/09 02:14 dwang rebooting target IOC
  • 290782 09/21/09 01:48 jpchen target motion/ioc reboot
  • 290780 09/21/09 01:41 jpchen target heater alarm and PID setpoint
  • 290776 09/21/09 01:07 dwang target alarms whenever beam trips
  • 290766 09/21/09 00:02 LeRose Shift Summary
  • Sunday

  • 290762 09/20/09 23:22 R. Subedi Spot check
  • 290759 09/20/09 23:00 R. Subedi Run after pvdis tagger in counting house restoration
  • 290758 09/20/09 22:42 R. Subedi pvdis gate generator in counting house middle room
  • 290755 09/20/09 22:17 Abdurahim Target strip charts
  • 290754 09/20/09 22:16 Abdurahim Target screen shots
  • 290753 09/20/09 22:12 Abdurahim Target IOC has been rebooted
  • 290750 09/20/09 21:30 R. Subedi Spot++ after target IOC reboot
  • 290741 09/20/09 20:28 LeRose No Beam
  • 290739 09/20/09 20:19 R. Subedi CODA soft reboot
  • 290734 09/20/09 18:37 R. Subedi Spot++
  • 290729 09/20/09 18:11 R. Subedi HapNoQweak to Happex
  • 290717 09/20/09 15:52 Bellini Reducing 02 % JT valve for L3
  • 290716 09/20/09 15:23 B.Quinn spot ++
  • 290711 09/20/09 14:26 B. Quinn roc31
  • 290707 09/20/09 14:12 mcnulty Here are some Qsq results from the Rarm for Hydrogen and Dummy Al
  • 290702 09/20/09 13:07 A. Camsonne ETROC stuck
  • 290701 09/20/09 12:58 mcnulty run record of beam position scan (with basic conclusions)
  • 290691 09/20/09 12:07 silwal spot++, -4mm in xi n BPM4B, walls hit
  • 290689 09/20/09 12:05 silwal spot++, -2mm in x in BPM4B, no walls hit yet
  • 290685 09/20/09 11:51 silwal spot ++, -5.5 mm in y in BPM4B, no walls hit yet
  • 290682 09/20/09 11:43 silwal spot++, -4mm in y in BPM4B, no walls hit yet
  • 290679 09/20/09 11:36 silwal spot++, -2mm in y in BPM4B, no walls hit yet
  • 290674 09/20/09 11:24 silwal spot++, +4 mm in y in BPM4B, no walls hit yet
  • 290671 09/20/09 11:12 silwal spot++, +2mm in y in BPM4B
  • 290667 09/20/09 11:01 silwal spot++, spot +4mm in x in BPM4B, wall hit
  • 290666 09/20/09 10:55 silwal spot++, +2mm in x in BPMB
  • 290665 09/20/09 10:54 A. Camsonne Compton low current run with coincidence 19543
  • 290663 09/20/09 10:50 A. Giusa Target stabilized
  • 290658 09/20/09 10:42 silwal spot++, beam @ the nominal center of the target
  • 290655 09/20/09 10:31 mcnulty s0 HV on, Beam current at 10uA (Preparing for beam pos scan)
  • 290653 09/20/09 10:15 mcnulty Beam position scan with smpling HRS DAQ only
  • 290650 09/20/09 09:23 Sarah Position Scan Run Plan II
  • 290649 09/20/09 09:19 bellini Actual target parameters after IOC1 reboot
  • 290648 09/20/09 09:16 bellini rebooting IOC13
  • 290645 09/20/09 08:49 bellini Regular behaviour on L3 target
  • 290644 09/20/09 08:35 Bellini Transient Red Alarm on target L3
  • 290642 09/20/09 08:31 Bellini Transient Red Alarm for Target L3
  • 290641 09/20/09 08:29 Bellini Transient Red Alarm for Target L3
  • 290635 09/20/09 08:20 apar Halfwave plate IN, Slug 10
  • 290631 09/20/09 07:58 mercado Owl Shift Summary
  • 290629 09/20/09 07:28 Yoomin Compton CODA rebooted
  • 290628 09/20/09 07:20 mercado Beam Energy Monitor
  • 290625 09/20/09 07:09 Kai target snapshot
  • 290624 09/20/09 06:21 Yoomin Compton CODA rebooted
  • 290612 09/20/09 03:48 Kai Pan target alarmed and recovered again
  • 290609 09/20/09 03:23 Kai target alarm
  • 290608 09/20/09 02:31 Yoomin CODA restarted
  • 290604 09/20/09 02:11 Yoomin Compton CODA reboot
  • 290601 09/20/09 00:55 Yoomin adaql2 screen2 monitor
  • 290595 09/20/09 00:24 Sarah Run Coordinator's Report for September 18 Swing to September 19 Day
  • 290592 09/20/09 00:07 Ramesh Subed Shift Summary
  • Saturday

  • 290590 09/19/09 23:58 R. Michaels Compton: back to production, test stopped
  • 290585 09/19/09 23:09 R. Michaels CompSimple run 19532
  • 290584 09/19/09 22:53 R. Subedi Stable beam
  • 290583 09/19/09 22:51 M Friend Compton RunControl Moved
  • 290582 09/19/09 22:48 R. Subedi Beam is quite stable for last two hours
  • 290581 09/19/09 22:44 R. Subedi Hall A Tools for check-list
  • 290576 09/19/09 20:30 M Friend spot++
  • 290564 09/19/09 19:24 Abdurahim Target strip charts
  • 290563 09/19/09 19:23 Abdurahim Target screen snapshot
  • 290560 09/19/09 17:57 R. Subedi Beam back
  • 290559 09/19/09 17:41 R. Subedi Beam away for sometime
  • 290556 09/19/09 17:13 A. Camsonne Resuming Compton production run 19529
  • 290555 09/19/09 17:04 Sarah Run Plan for Tonight and Tomorrow
  • 290554 09/19/09 17:00 M Friend spot++ with beam!
  • 290550 09/19/09 16:49 M Friend spot ++
  • 290546 09/19/09 16:46 M Friend Wrong Pedestal Value?
  • 290545 09/19/09 16:41 A. Camsonne EtrocFifo dma configuration
  • 290539 09/19/09 16:10 silwal PC IHWP IN/OUT voltages mixup, and corrected
  • 290538 09/19/09 15:57 Nuruzzaman Day Shift Summary
  • 290537 09/19/09 15:56 A. Camsonne Compton stand alone run 19520 ETROCFifo
  • 290535 09/19/09 15:53 A. Giusa grab of target screens
  • 290534 09/19/09 15:50 A. Camsonne Compton stand alone run 19519 ETROCOnly
  • 290533 09/19/09 15:50 R. Michaels new CompSimple config
  • 290530 09/19/09 15:47 A. Camsonne Compton stand alone run 19518 Test Dma ETROCFifo
  • 290529 09/19/09 14:53 Nuruzzaman spot++
  • 290515 09/19/09 12:08 Nuruzzaman CODA restarted, changed config. to HapNoQweak
  • 290504 09/19/09 09:49 Nuruzzaman Compton CODA restarted,Run 19514 started
  • 290501 09/19/09 08:52 A. Giusa Grab of target screens
  • 290494 09/19/09 08:01 mercado Owl Shift Summary
  • 290491 09/19/09 07:40 mercado Beam Energy Monitor
  • 290488 09/19/09 06:58 R. Michaels stop compton for halfwave flip
  • 290487 09/19/09 06:58 apar Halfwave plate OUT, Slug 9
  • 290486 09/19/09 06:53 Yoomin spot++ of run 24766
  • 290483 09/19/09 06:51 mercado Halfwave plate change to OUT
  • 290481 09/19/09 06:49 apar Halfwave plate OUT, Slug 9
  • 290477 09/19/09 06:32 Yoomin Target Screen capture
  • 290467 09/19/09 03:55 R. Michaels proposed simplified Compton DAQ
  • 290466 09/19/09 03:49 R. Michaels Plane D asymmetries too
  • 290465 09/19/09 03:47 R. Michaels Compton Edet Asymmetries
  • 290460 09/19/09 02:24 R. Michaels Compton to production mode
  • 290459 09/19/09 01:55 R. Michaels EtrocOnly runs 19508 and 9
  • 290456 09/19/09 01:50 R. Michaels Compton Edet hit spectrum
  • 290455 09/19/09 01:46 mercado Back to HAPPEX config
  • 290452 09/19/09 01:19 R. Michaels Compton HV trip
  • 290451 09/19/09 00:51 R. Michaels cavity flip started
  • 290448 09/19/09 00:41 R. Michaels More EtrocOnly running
  • 290447 09/19/09 00:25 paschke Beam energy monitor snapshot
  • 290441 09/19/09 00:14 R. Michaels EtrocOnly config run
  • Friday

  • 290439 09/18/09 23:59 paschke Swing Shift Summary
  • 290437 09/18/09 23:46 M Friend spot++
  • 290434 09/18/09 23:42 paschke access to reset moller quads
  • 290430 09/18/09 22:59 A. Camsonne Standard ETROC firmware loaded with photon input on pu_in[7]
  • 290429 09/18/09 22:33 paschke expert coming in for access on moller quads
  • 290428 09/18/09 22:21 paschke Moller quads not resetting
  • 290427 09/18/09 22:15 paschke Moller quads trip
  • 290426 09/18/09 22:12 A. Camsonne ETROC photon trigger using pu_in[7] input
  • 290425 09/18/09 22:09 M Friend spot++
  • 290418 09/18/09 22:00 A. Camsonne ETROC photon trigger only firmware loaded
  • 290415 09/18/09 20:54 A. Camsonne Compton disk cleanup continued
  • 290409 09/18/09 20:12 M Friend spot++
  • 290404 09/18/09 20:05 A. Camsonne Compton ROC0 error message
  • 290403 09/18/09 19:41 A. Camsonne Compton ROC6 error message
  • 290402 09/18/09 19:37 Sarah Run Coordinator's Report for September 17 Swing to September 18 Day
  • 290400 09/18/09 18:52 Sarah Run Plan for Tonight
  • 290397 09/18/09 18:16 M Friend Compton comptuter crashed
  • 290396 09/18/09 17:33 M Friend spot++
  • 290389 09/18/09 16:34 A. Camsonne High statistics run 19482
  • 290383 09/18/09 16:19 A. Camsonne Compton run 19488
  • 290382 09/18/09 16:15 Bellini Target snapshot
  • 290381 09/18/09 16:14 A. Camsonne Compton / adaq disk clean up
  • 290378 09/18/09 16:00 Nuruzzaman Day Shift Summary
  • 290377 09/18/09 15:55 A. Giusa target screens referred by my previous entry
  • 290375 09/18/09 15:52 A. Giusa grab of target screens
  • 290374 09/18/09 15:43 A. Camsonne Compton DAQ stable with 10 % prescale
  • 290372 09/18/09 15:15 Nuruzzaman spot++
  • 290371 09/18/09 15:13 A. Camsonne Error message in coincidence mode
  • 290367 09/18/09 14:21 A. Camsonne Compton electron detector threshold scan
  • 290366 09/18/09 14:18 A. Giusa heater power
  • 290365 09/18/09 14:16 jpchen HP heater problem
  • 290362 09/18/09 13:43 R. Subedi HV for PVDIS bar need to be always ON
  • 290359 09/18/09 13:14 A. Camsonne Compton electron detector multiplicity set to 2 planes threshold 11 for un 19481
  • 290358 09/18/09 13:08 A. Camsone Compton electron detector prompt analysis Run 19480
  • 290357 09/18/09 12:54 A. Camsonne Prompt analysis Compton 19479
  • 290356 09/18/09 12:52 A. Camsonne Compton electron detector moved to position 1mm
  • 290355 09/18/09 12:38 A. Camsonne Compton rates dead time effect
  • 290354 09/18/09 12:31 M Friend Compton FADC config chang
  • 290353 09/18/09 12:24 Nuruzzaman spot++ snaps
  • 290337 09/18/09 11:54 R. Michaels cron and epics playing nicely now
  • 290335 09/18/09 11:53 A. Camsonne Compton electron detector in position 4 mm
  • 290333 09/18/09 11:47 A. Giusa target temperature and beam trips
  • 290332 09/18/09 11:41 R. Michaels test compton getscaler
  • 290331 09/18/09 11:32 R. Michaels compton getscaler in cron
  • 290330 09/18/09 11:01 Nuruzzaman CODA configuration changed to HapNoQweak
  • 290322 09/18/09 09:32 A. Giusa cryoflows - JT valve
  • 290321 09/18/09 09:10 Kowalski High Voltage
  • 290320 09/18/09 08:42 Nuruzzaman CODA restarted
  • 290310 09/18/09 08:04 Ahmed. ABU and BANU report.
  • 290309 09/18/09 07:54 adaq spot++
  • 290308 09/18/09 07:50 ZAFAR AHMED Owl shift summary
  • 290305 09/18/09 07:04 deng target JT valve changed from 54% to 53.5%
  • 290304 09/18/09 06:45 paschke compton daq: scrolling error, ended run #19473
  • 290301 09/18/09 06:35 Ahmed Spot++
  • 290300 09/18/09 05:42 deng target: JT valve changed from 54.0% to 54.5% then back to 54.0%
  • 290297 09/18/09 05:16 silwal Beam Position scan
  • 290296 09/18/09 05:14 deng target: increase JT value from 53.5% to 54.0%
  • 290295 09/18/09 04:59 silwal Beam Position scan
  • 290286 09/18/09 04:07 Ahmed. spot ++ Hydrogen target. Beam was moved to -2mm in Y direction at 4B.
  • 290283 09/18/09 04:03 deng target moved to LH2 from holy carbon
  • 290282 09/18/09 04:02 Ahmed. spot++. Carbon holy target. Beam moved to -2mm in Y direction at 4B.
  • 290279 09/18/09 03:50 Ahmed. spot++. Carbon holy target. Position(0,1) at 4B.
  • 290276 09/18/09 03:44 deng target moved to holy carbon
  • 290274 09/18/09 03:20 Ahmed. spot++. Hydrogen target. Position (0,1) at 4B.
  • 290269 09/18/09 02:43 Ahmed. spot++ for hydrogen target. Beam is moved to -2mm in X at 4B.
  • 290265 09/18/09 02:37 Ahmed. spot++ for holy carbon. Beam i s moved to -2mm in X at 4B.
  • 290264 09/18/09 02:36 deng target moved to LH2 from holy carbon
  • 290261 09/18/09 02:25 Ahmed. spot++ holy carbon +1mm in X for 4B
  • 290258 09/18/09 02:17 deng target moved to carbon hole
  • 290254 09/18/09 01:50 Ahmed. spot++ beam moved +1mm BPM 4b
  • 290251 09/18/09 01:37 deng snap shot of target tool--heater strip
  • 290250 09/18/09 01:35 deng snap shot of tartget tools
  • 290249 09/18/09 01:34 Ahmed. BPM strips snap shot.
  • 290248 09/18/09 01:33 Ahmed Hall A tools.
  • 290246 09/18/09 01:13 Ahmed. spot++ hydrogen target.
  • 290243 09/18/09 01:06 deng target moved to LH2 from carbon hole
  • 290242 09/18/09 01:03 deng target moved to carbon hole at 00:45 am
  • 290241 09/18/09 01:01 Ahmed spot++ LHRS run 24746
  • 290238 09/18/09 00:53 ZAFAR AHMED spot++ for holy carbon 5uA.
  • 290229 09/18/09 00:20 ZAFAR AHMED spot++
  • 290228 09/18/09 00:02 jpchen swing shift summary
  • 290226 09/18/09 00:01 Xinhu Yan Spot++
  • Thursday

  • 290223 09/17/09 23:57 M Friend Good Compton Rates
  • 290220 09/17/09 23:32 jpchen bpm/bcm charts
  • 290219 09/17/09 23:30 jpchen tools screen
  • 290218 09/17/09 23:30 M Friend Compton Beam Position
  • 290217 09/17/09 23:30 jpchen Compton chart
  • 290216 09/17/09 23:11 silwal IHWP changed to IN, flipper did not work, but slug number incremented by 1
  • 290214 09/17/09 23:01 jpchen beam half-wave plate IN
  • 290210 09/17/09 21:55 adaq Hall-A tools screen
  • 290207 09/17/09 21:37 A. Camsonne Edet moved to 7 mm
  • 290206 09/17/09 21:35 Sarah Run Coordinator's Report for September 16 Swing to September 17 Day
  • 290205 09/17/09 21:28 Sarah Temperature in Counting House
  • 290204 09/17/09 21:24 jpchen Compton HV trip, reset.
  • 290203 09/17/09 21:01 Sarah Position Scan Run Plan
  • 290202 09/17/09 20:56 Bo Target main screen
  • 290201 09/17/09 20:56 A. Camsonne Compton electron detector in position 3 mm
  • 290200 09/17/09 20:54 Bo Target strip charts
  • 290199 09/17/09 20:52 R. Michaels scaler fixed (maybe !!)
  • 290189 09/17/09 20:23 Bo cryoflow adjustment
  • 290183 09/17/09 20:20 Xinhu Yan Spot++
  • 290180 09/17/09 20:11 silwal for bad runs, bcm10 stays bad for the entire length of the run
  • 290177 09/17/09 19:35 R. Michaels CH scaler (again)
  • 290176 09/17/09 19:32 adaq compton DAQ hung up, restarted
  • 290175 09/17/09 19:27 jpchen Compton
  • 290172 09/17/09 18:32 Sarah Run Plan for Tonight
  • 290171 09/17/09 18:29 M Friend Compton Data
  • 290170 09/17/09 18:18 jpchen adjusted cryoflow to have ~50 watts with beam on
  • 290169 09/17/09 18:10 jpchen HP heater not responding, rebooted ioc
  • 290168 09/17/09 18:01 Xinhu Yan Blumi sum asym online analysis
  • 290167 09/17/09 17:57 silwal CH scaler sync bad again, Hall C cm readout is bad
  • 290164 09/17/09 17:32 jpchen Compton running
  • 290163 09/17/09 17:29 jpchen beam checked, production resumed with 98 uA.
  • 290162 09/17/09 17:27 Xinhu Yan BLumi width of asym peak
  • 290161 09/17/09 17:17 A. Camsonne Compton rates good
  • 290159 09/17/09 16:57 jpchen harp scan results
  • 290158 09/17/09 16:49 Xinhu Yan Spot++
  • 290144 09/17/09 16:02 Holmstrom Day Shift Summary
  • 290142 09/17/09 15:59 Nuruzzaman Injector Vs Synchronization check: 5 more test runs taken
  • 290141 09/17/09 15:57 O.Glamazdin Moller results
  • 290131 09/17/09 15:38 A. Giusa Grab of target screens
  • 290130 09/17/09 15:36 Holmstrom Hall A Tools
  • 290129 09/17/09 15:31 Nuruzzaman Injector Vs LHRS Synchronization plots looks good now
  • 290113 09/17/09 15:03 R. Michaels CH scaler replaced
  • 290108 09/17/09 14:58 M Friend Compton Data
  • 290104 09/17/09 14:14 A. Camsonne Compton Electron detector out
  • 290098 09/17/09 12:16 A. Giusa Target moved to empty
  • 290097 09/17/09 12:12 A. Camsonne Compton vertical scan
  • 290096 09/17/09 12:07 A. Giusa target temperature and beam trip
  • 290094 09/17/09 12:01 A. Camsonne Compton attenuation gain set to 6.2
  • 290091 09/17/09 11:38 A. Camsonne Compton Electron detector in beam position 4mm
  • 290090 09/17/09 11:05 A. Giusa target freezing
  • 290089 09/17/09 10:49 Nuruzzaman BCM signals looking different.
  • 290086 09/17/09 10:03 R. Michaels suggestion re: CH scaler
  • 290085 09/17/09 09:55 silwal Scaler state for hall C bcm readout good/bad
  • 290084 09/17/09 09:51 silwal Injector Happex Scaler Sync
  • 290081 09/17/09 09:35 silwal mistake on CH scaler sync earlier, when hall C bcm is not read out properly, the CH scaler sync is bad
  • 290080 09/17/09 09:13 A. Giusa Grab of target screens
  • 290079 09/17/09 09:00 silwal CH scaler sync monitors fine, even for the run when Hall C bcm was not being read out properly
  • 290078 09/17/09 08:59 Nuruzzaman raster looks fine
  • 290077 09/17/09 08:53 silwal went back to the old crl, CH scaler sync off again
  • 290067 09/17/09 08:35 silwal Injector Happex scaler out of synch with modified crl, intended to fix bcm10 readout.
  • 290058 09/17/09 08:06 silwal target screen
  • 290057 09/17/09 08:01 silwal owl shift summary
  • 290055 09/17/09 07:45 R. Michaels re: CH scaler getting stuck
  • 290054 09/17/09 07:21 Yoomin ROC31 hung up. CODA reset
  • 290051 09/17/09 07:05 silwal Injector BPM differences, rms as big as 58 um!
  • 290048 09/17/09 06:44 silwal beam back @ 100 uA since 6:45 am
  • 290047 09/17/09 06:38 silwal beam down ~30 minutes, ioc down, is already rebooted, should be back by 7:00 am
  • 290044 09/17/09 05:16 silwal Hall C bcm readout back to normal
  • 290041 09/17/09 04:51 silwal beam energy monitor snapshot
  • 290035 09/17/09 04:19 silwal Hall C bcm readout died, CH scaler out of sync
  • 290030 09/17/09 03:09 mcnulty Target screens
  • 290029 09/17/09 02:20 Yoomin spot++
  • 290022 09/17/09 00:34 M Friend Compton's looking good
  • 290013 09/17/09 00:00 Ole Hansen Swing Shift summary
  • Wednesday

  • 290012 09/16/09 23:37 gjin another trip of H2 gas panel fill side pressure
  • 290011 09/16/09 23:27 Ole Hansen Beam is back, resuming production
  • 290009 09/16/09 23:16 gjin target tools screen snap shot
  • 290008 09/16/09 23:09 gjin H2 fill pressure out of lower range for an instant
  • 290007 09/16/09 23:02 Ole Hansen Beam positions off
  • 290006 09/16/09 22:58 Ole Hansen spot++
  • 290002 09/16/09 22:49 Ole Hansen Resuming production
  • 290001 09/16/09 22:48 Ole Hansen Larger instrinsic beam spot size
  • 290000 09/16/09 22:46 gjin target->H2
  • 289999 09/16/09 22:24 Ole Hansen Spot size
  • 289998 09/16/09 22:23 Ole Hansen Data taking
  • 289996 09/16/09 22:00 gjin target->H2
  • 289995 09/16/09 22:00 Ole Hansen Production
  • 289993 09/16/09 21:49 Ole Hansen spot++
  • 289991 09/16/09 21:42 Ole Hansen Harp results
  • 289990 09/16/09 21:02 Ole Hansen Beam ready
  • 289984 09/16/09 19:52 Ole Hansen We have beam
  • 289983 09/16/09 18:22 Sarah Run Coordinator's Report for September 15 Swing to September 16 Day
  • 289981 09/16/09 17:37 Sarah The Run Plan for Tonight and Tomorrow
  • 289980 09/16/09 17:00 mercado Back to delayed helicity
  • 289974 09/16/09 16:14 gomez iocha14/iocha16 back on hlasrv
  • 289972 09/16/09 16:04 Kalyan Shift summary / Run summary
  • 289971 09/16/09 16:03 Kalyan R-HRS Q2 is back - remote reset did it.
  • 289968 09/16/09 15:46 Kalyan Right arm Q2 quenched
  • 289967 09/16/09 15:04 mercado Helicity switched to 'In-Time' with no delay
  • 289964 09/16/09 14:53 A. Giusa Target moved to empty
  • 289963 09/16/09 14:26 adaq beam off
  • 289962 09/16/09 14:14 Kalyan update on beam
  • 289959 09/16/09 13:28 Kalyan beam status
  • 289958 09/16/09 13:19 silwal back to "Happex" CODA config
  • 289954 09/16/09 12:46 R. Michaels possible CH scaler fix
  • 289953 09/16/09 12:40 dalton Slug number really 7
  • 289952 09/16/09 12:38 Kalyan beam positions not good
  • 289944 09/16/09 12:23 A. Giusa target moved to carbon.
  • 289934 09/16/09 11:53 Kalyan Beam off for compton tune
  • 289930 09/16/09 11:48 Kalyan Charge feedback found to be off - turned on now.
  • 289929 09/16/09 11:45 R. Michaels CH scaler problem
  • 289927 09/16/09 11:19 Kalyan Production with 100uA
  • 289925 09/16/09 11:18 M Friend Root version change
  • 289923 09/16/09 11:16 R. Michaels spot_L repaired, present picture
  • 289919 09/16/09 10:48 Kalyan Hall-A dpp value dropped
  • 289913 09/16/09 10:31 apar Halfwave plate OUT, Slug 11
  • 289912 09/16/09 10:31 apar Halfwave plate OUT, Slug 10
  • 289911 09/16/09 10:29 apar Halfwave plate OUT, Slug 9
  • 289910 09/16/09 10:28 apar Halfwave plate OUT, Slug 8
  • 289908 09/16/09 10:22 R. Michaels Mon saturation transient
  • 289907 09/16/09 10:18 Nuruzzaman BPM 12X is saturating
  • 289906 09/16/09 10:13 Kalyan Hall-A tools and BPM strip charts
  • 289905 09/16/09 10:02 R. Michaels production starting
  • 289903 09/16/09 09:57 giusa target moved to H2
  • 289902 09/16/09 09:55 R. Michaels spot check -- good -- ready for production
  • 289900 09/16/09 09:45 gomez iocha16 (right HRS magnet control) failure to reboot
  • 289898 09/16/09 09:38 V. Sulkosky L-HRS Spot++ Problem
  • 289896 09/16/09 09:20 mcnulty going to production mode...setting HV's
  • 289895 09/16/09 09:17 mcnulty VDC's back to zero HV
  • 289894 09/16/09 09:16 mcnulty Work done during access/ spot++ problem investigation
  • 289876 09/16/09 08:19 Kalyan VDCs on both arms turned on
  • 289872 09/16/09 07:52 mcnulty spot++ trouble/ no data seen in Larm
  • 289859 09/16/09 07:00 mcnulty MCC doing more tuning in hall A line
  • 289857 09/16/09 06:53 mcnulty spot++ errors! #@%$#@
  • 289854 09/16/09 06:51 deng snap shot of target tools
  • 289850 09/16/09 06:42 deng target moved to carbon hole
  • 289849 09/16/09 06:41 mcnulty MCC harp scan software not working
  • 289847 09/16/09 06:26 mcnulty MCC tuning beam into hall now
  • 289846 09/16/09 06:20 R. Michaels re: CH scaler
  • 289843 09/16/09 05:45 mcnulty Going to Beam Permit
  • 289828 09/16/09 00:22 mcnulty update on beam delivery status
  • Tuesday

  • 289821 09/15/09 23:56 hoyoung Swing shift summary
  • 289820 09/15/09 23:48 Kai Access Activity (PVDIS, Kai Pan)
  • 289819 09/15/09 23:20 R. Michaels re: wierd spot_R
  • 289817 09/15/09 23:10 R. Michaels CH scaler issues
  • 289816 09/15/09 23:01 silwal CH scaler working fine
  • 289809 09/15/09 22:21 R. Subedi Access activity (pvdis work, late entry)
  • 289796 09/15/09 21:24 gjin target tools screen
  • 289795 09/15/09 21:23 gjin target screen snap shot
  • 289794 09/15/09 21:05 Sarah Run Coordinator's Report for September 14 Swing to September 15 Da
  • 289793 09/15/09 20:58 silwal backed out of injector studies setup
  • 289790 09/15/09 20:35 dalton, silw Injector studies
  • 289780 09/15/09 19:58 M Friend Compton photon detector moved
  • 289776 09/15/09 19:43 Sarah Run Plan for Tonight
  • 289774 09/15/09 19:35 Sarah Accessed the Hall Again to Close the Spectrometer Doors
  • 289771 09/15/09 18:45 deng access activity ( pvdis DAQ work)
  • 289770 09/15/09 18:45 Sarah No Right Arm Tonight
  • 289764 09/15/09 18:22 folts tech magnets
  • 289761 09/15/09 18:07 A. Camsonne CVS update of Compton software acpa
  • 289756 09/15/09 17:10 Sarah To All Shift Takers in Hall A
  • 289751 09/15/09 16:21 Bob/Alex Compton tried/failed to use new ROC1/2 cpus
  • 289750 09/15/09 16:21 jpchen Day shift summary
  • 289741 09/15/09 15:58 mercado Run 12347 is Pedestal Run with In-time Helicity
  • 289738 09/15/09 15:15 jpchen target moved to EMPTY, target screen
  • 289737 09/15/09 15:14 jpchen target charts
  • 289723 09/15/09 13:46 reimer Target moved to optics
  • 289717 09/15/09 13:27 reimer IOCHA16 reboot
  • 289694 09/15/09 12:22 reimer IOCHA16 reboot
  • 289686 09/15/09 12:08 M Friend Epics Reboot
  • 289683 09/15/09 11:12 reimer Smoke? & water
  • 289682 09/15/09 11:06 M Friend Scaler channels added to FADC DAQ
  • 289679 09/15/09 10:55 M Friend Multiple Compton HV Trips
  • 289676 09/15/09 10:27 R. Subedi Left HRS vdc and S0 turned OFF
  • 289653 09/15/09 09:48 R. Subedi left HRS vdc and S0 turned ON
  • 289644 09/15/09 08:26 paschke injector studies
  • 289632 09/15/09 07:46 ZAFAR AHMED Hall A open, Controlled access.
  • 289628 09/15/09 07:39 ZAFAR AHMED Owl shift summary.
  • 289625 09/15/09 07:32 paschke Dithering OFF. Beam studies at MCC. Parity DAQ in use.
  • 289622 09/15/09 05:58 ZAFAR AHMED Lecroy HV
  • 289621 09/15/09 05:50 ZAFAR AHMED Hall A tools
  • 289619 09/15/09 05:35 Yoomin spot++ after beam back to 100uA
  • 289616 09/15/09 05:16 Jack (Techs) Right Q1
  • 289615 09/15/09 05:01 ZAFAR AHMED Controlled access.
  • 289608 09/15/09 04:09 A. Camsonne Compton DAQ up, asked to MCC for tuning to reduce rate
  • 289607 09/15/09 04:03 Yoomin CODA restarted
  • 289605 09/15/09 03:58 A. Camsonne Compton detectors moved out
  • 289600 09/15/09 03:47 Antonio Gius grab of running things
  • 289599 09/15/09 03:42 ZAFAR AHMED Spot++
  • 289598 09/15/09 03:09 giusa Power to Zero
  • 289595 09/15/09 02:53 Yoomin Spot++
  • 289592 09/15/09 02:38 ZAFAR AHMED 66uA Current.
  • 289591 09/15/09 02:26 ZAFAR AHMED Right arm Q1 is down.
  • 289590 09/15/09 02:07 dalton bpm12 can exceed 13200 without saturating
  • 289584 09/15/09 00:07 adaq HWP changed to "OUT"
  • 289583 09/15/09 00:06 hoyoung Swing shift summary
  • 289581 09/15/09 00:04 apar Halfwave plate OUT, Slug 7
  • Monday

  • 289577 09/14/09 23:16 silwal Hall C bcm moved to bcm10 from bcm6 for historical reasons
  • 289576 09/14/09 23:15 silwal the defunct spot++ was from right HRS
  • 289575 09/14/09 23:15 silwal the defunct spot++ was from right HRS
  • 289574 09/14/09 23:01 silwal spot++ ok on the left arm
  • 289569 09/14/09 22:47 silwal weird spot++, nothing on the bpmA
  • 289555 09/14/09 21:17 A. Camsonne Compton counting photon attenuator DAQ lowered to 3
  • 289552 09/14/09 20:07 reimer target alarms
  • 289551 09/14/09 19:59 silwal Hall C bcm signal: bcm6 on datamap
  • 289550 09/14/09 19:53 Sarah Hall A Access Activities for September 15, 2009
  • 289549 09/14/09 19:42 silwal parity DAQ config changed from HapNoQweak to Happex, db files unchanged
  • 289547 09/14/09 19:32 mcnulty BCM signal from hall C to hall A has been connected
  • 289546 09/14/09 19:02 Sarah Run Coordinator's Report for September 13 Swing to September 14 Day
  • 289545 09/14/09 18:44 M Friend Good Compton Photon Rates
  • 289542 09/14/09 17:49 adaq spot++
  • 289540 09/14/09 17:33 Sarah Run Plan for Tonight
  • 289539 09/14/09 16:58 hoyoung Beam trip
  • 289537 09/14/09 16:35 hoyoung Compton HV reset
  • 289535 09/14/09 16:25 jpchen day shift summary
  • 289533 09/14/09 16:18 reimer Target Strip Charts
  • 289532 09/14/09 16:15 reimer target display
  • 289531 09/14/09 15:52 jpchen target screen
  • 289530 09/14/09 15:52 jpchen target charts
  • 289529 09/14/09 15:36 fansler LQ1 Momentum
  • 289528 09/14/09 15:29 Jack (Techs) LQ1
  • 289521 09/14/09 14:42 Sarah Controlled Access to Reset Q1
  • 289520 09/14/09 14:06 A. Camsonne Compton photon electron coincidence run 19363
  • 289519 09/14/09 14:02 Sarah Q1 is behaving strangely
  • 289517 09/14/09 13:59 M Friend Compton Rates
  • 289512 09/14/09 12:59 A. Camsonne Run19355 standalone position 2mm
  • 289511 09/14/09 12:37 A. Camsonne ETROC in beam at position 2mm
  • 289508 09/14/09 12:16 A. Camsonne Junk runs until run 19362
  • 289507 09/14/09 12:01 A. Camsonne ETROC stand alone run 19355
  • 289506 09/14/09 11:42 jpchen beam back to orginal configuration before tuning
  • 289503 09/14/09 10:53 A. Camsonne Compton electron photon timing
  • 289502 09/14/09 10:48 M Friend Compton Background Rates!!!
  • 289501 09/14/09 10:40 jpchen exit analyzer properly by .qqqqqqqq (8 q's)
  • 289500 09/14/09 10:35 jpchen Bema tuning not good for Compton. Need re-tuning.
  • 289498 09/14/09 10:30 jpchen HVs
  • 289497 09/14/09 10:19 jpchen beam back
  • 289495 09/14/09 10:13 jpchen beam tuning since 9:45
  • 289491 09/14/09 09:23 adaq Spot++
  • 289488 09/14/09 08:17 Xinhu Yan Target Strip Charts
  • 289487 09/14/09 07:59 ZAFAR AHMED OWL SHIFT SUMMARY
  • 289486 09/14/09 07:53 deng snapshot of target tools
  • 289477 09/14/09 06:37 ZAFAR AHMED No beam for next 2 hours.
  • 289473 09/14/09 05:10 A. Camsonne Beamline HV connection lost
  • 289470 09/14/09 04:45 ZAFAR AHMED Hall A tools.
  • 289463 09/14/09 01:08 A. Camsonne Compton ru 19333 Stand alone electron
  • 289460 09/14/09 00:32 A. Camsonne Cavity not locking / EPICS IOC reboot
  • 289459 09/14/09 00:03 LeRose Shift Summary
  • Sunday

  • 289455 09/13/09 23:17 Armstrong prompt analysis issue
  • 289454 09/13/09 23:04 kai spot++ #24707
  • 289452 09/13/09 22:57 LeRose Beam Back
  • 289448 09/13/09 22:50 LeRose No Beam
  • 289440 09/13/09 22:38 Armstrong some junk runs
  • 289413 09/13/09 22:04 paschke dithering orthogonality still tenuous
  • 289412 09/13/09 21:42 LeRose Hall A Tools/Beam
  • 289409 09/13/09 21:18 Kai reboot ROC of TS on Left-arm. Left coda works now.
  • 289408 09/13/09 21:17 Armstrong Dithering looks singular
  • 289407 09/13/09 21:08 paschke helicity pulled from Counting House. Only in injector now!
  • 289406 09/13/09 21:03 Kai snapshot of target tools
  • 289401 09/13/09 20:39 Armstrong low beam cut readjusted for 100 muA
  • 289397 09/13/09 20:31 LeRose Beam Back
  • 289394 09/13/09 20:25 Kai spot++ on LH2, #4569
  • 289392 09/13/09 20:17 Kai spot++ study
  • 289379 09/13/09 19:35 M Friend Compton signal/background improvement
  • 289374 09/13/09 18:15 A. Camsonne Compton single spectrum
  • 289372 09/13/09 18:09 LeRose No Beam
  • 289370 09/13/09 17:53 Sarah Run Coordinator's Report for September 12 Swing to September 13 Day
  • 289369 09/13/09 17:52 Sarah Run Plan for Tonight
  • 289366 09/13/09 17:08 paschke Low beam cut lowered for 80uA running
  • 289362 09/13/09 16:31 A. Camsonne Compton Electron detector in beam position 3mm
  • 289361 09/13/09 16:04 jroche day shift summary
  • 289360 09/13/09 15:48 mercado Pedestal Means for InTime/Delayed Helicity... Gaussian Fits
  • 289357 09/13/09 15:29 jroche magic?
  • 289355 09/13/09 15:16 jroche slowly ramping up the beam
  • 289354 09/13/09 15:09 mercado Pedestal Diff Mean Trend - Better image
  • 289353 09/13/09 14:58 mercado Comparing Pedestals between In-time/Delayed helicity
  • 289352 09/13/09 14:45 M Friend Compton HV
  • 289350 09/13/09 14:00 M Friend Config back to normal
  • 289349 09/13/09 13:47 A. Camsonne HV Compton HV power supply specs
  • 289346 09/13/09 13:26 jroche lots of beam trip
  • 289345 09/13/09 13:12 M Friend Compton run config change
  • 289344 09/13/09 13:02 mercado BCM Unser now has ~40,000 Signal
  • 289339 09/13/09 11:01 jroche data look good to me
  • 289336 09/13/09 10:07 mercado Target Screen
  • 289335 09/13/09 10:07 M Friend Compton Base Current Shifts
  • 289334 09/13/09 09:57 roche Hall A tools
  • 289332 09/13/09 09:49 jroche S0 HV off
  • 289331 09/13/09 09:38 Kalyan beam is back - spot++
  • 289328 09/13/09 08:47 jroche No beam for a while
  • 289327 09/13/09 08:09 V. Sulkosky Owl Shift Summary
  • 289322 09/13/09 07:47 A. Camsonne Compton photon HV current steady at 5 uA
  • 289317 09/13/09 07:32 Yoomin Beam check
  • 289315 09/13/09 07:16 Xinhu Yan Target Charts
  • 289314 09/13/09 07:02 A. Camsonne Compton photon rates look good - increasing HV to 1875 V
  • 289313 09/13/09 06:59 V. Sulkosky Full Description of Beam Problem
  • 289312 09/13/09 06:35 A. Camsonne Compton electron and photon detector moved out for tune, laser off for background check
  • 289308 09/13/09 06:17 V. Sulkosky Tune Beam to Hall A
  • 289305 09/13/09 04:58 V. Sulkosky Beam Update (1 hour)
  • 289304 09/13/09 04:34 V. Sulkosky Beam Update (Maybe 2 hours)
  • 289301 09/13/09 03:48 V. Sulkosky Hall A Tools
  • 289300 09/13/09 03:43 V. Sulkosky Parity Runs 12260-12266 Are Cosmics
  • 289297 09/13/09 02:43 Yoomin Restart CODA
  • 289293 09/13/09 02:10 Xinhu Yan Screenshot for Target Strip Charts
  • 289292 09/13/09 01:45 V. Sulkosky Beam Update
  • 289289 09/13/09 00:33 V. Sulkosky Description of Beam Problems
  • 289286 09/13/09 00:14 Franklin Compton Accumulator Helicity Pair Plots
  • Saturday

  • 289284 09/12/09 23:54 jroche Swing shift summary
  • 289283 09/12/09 23:44 Kiadtisak Target Gui Screenshot
  • 289282 09/12/09 23:43 Kiadtisak Screenshot for Target Strip Charts
  • 289277 09/12/09 22:06 Kiadtisak Target moved to empty
  • 289276 09/12/09 21:54 jroche no beam
  • 289271 09/12/09 19:48 jroche beam "lost" in the injector
  • 289269 09/12/09 19:41 dalton IHWP changed to IN, slug 6
  • 289267 09/12/09 18:52 M Friend Laser Probably Not Having Trouble
  • 289266 09/12/09 18:44 M Friend Laser Having Trouble?
  • 289265 09/12/09 18:30 jroche hall a tool capture
  • 289262 09/12/09 17:58 Sarah Run Plan for Tonight
  • 289261 09/12/09 17:44 jroche current beam size (harps)
  • 289260 09/12/09 17:42 Sarah Run Coordinator's Report for September 11 Swing to September 12 Day Shift
  • 289258 09/12/09 17:23 Kiadtisak Target moved to LH2
  • 289257 09/12/09 17:16 A. Camsonne Compton electron detector moved to position 2mm
  • 289256 09/12/09 17:15 M Friend Compton HV
  • 289255 09/12/09 17:00 Kiadtisak target moved to empty
  • 289253 09/12/09 16:29 Kiadtisak Target Gui Screenshot
  • 289252 09/12/09 16:28 Kiadtisak Screenshot for Target Strip Charts
  • 289250 09/12/09 16:19 silwal monsatcut changed to 132000, bpm12 not saturated at 131000 yet
  • 289247 09/12/09 16:05 A. Deur Day shift summary
  • 289242 09/12/09 14:00 silwal 316x176um at the target, harp scan from 1:10 am, 09/12/09, Saturday
  • 289241 09/12/09 13:53 silwal 225x179 um spot size at the target, harp scan from 6:30 am, 09/11/09, Friday
  • 289240 09/12/09 13:53 silwal 220x170 um spot size at target, harp scan from 4:00 am, 09/11/09, Friday
  • 289239 09/12/09 13:30 A. Deur BEM snapshot
  • 289236 09/12/09 13:15 dalton Target screenshot
  • 289235 09/12/09 13:07 A. Deur BPM striptool snapshot
  • 289232 09/12/09 12:42 Nuruzzaman Raster looks fine
  • 289229 09/12/09 12:30 Jack (Techs) Right Q2
  • 289228 09/12/09 12:06 mercado Looking at boiling results
  • 289225 09/12/09 11:40 Nuruzzaman CODA restarted
  • 289224 09/12/09 11:32 R. Michaels raster X,Y flipped, I agree
  • 289217 09/12/09 09:50 dalton Target screen capture
  • 289216 09/12/09 09:21 A. Deur Hall A tools
  • 289215 09/12/09 09:07 Sarah Access into Hall A at 10:00am
  • 289214 09/12/09 09:07 dwang compton Edet efficiency
  • 289213 09/12/09 09:04 dwang/A.Cams compton spectrum
  • 289209 09/12/09 08:37 A. Deur Compton HV keeps tripping
  • 289208 09/12/09 08:19 V. Sulkosky Owl Shift Summary
  • 289207 09/12/09 08:18 V. Sulkosky re: raster X vs Y
  • 289203 09/12/09 07:09 V. Sulkosky Beam Energy Monitor
  • 289200 09/12/09 06:20 silwal boiling studies, owl shift
  • 289197 09/12/09 05:23 jroche Hall A tool capture
  • 289196 09/12/09 05:19 V. Sulkosky BPM FFB Gain in Manual for Production
  • 289195 09/12/09 05:10 silwal changed monsatcut to 200000
  • 289191 09/12/09 04:35 V. Sulkosky BPM FFB gain Control: Auto
  • 289190 09/12/09 04:34 Xinhu Yan Boiling test for LH2 target
  • 289188 09/12/09 03:38 V. Sulkosky Compton Rate Impact due to prebuncher adjustments
  • 289187 09/12/09 03:33 V. Sulkosky Full BCM Stripchart
  • 289186 09/12/09 03:26 V. Sulkosky BPM Gain control now Manual
  • 289184 09/12/09 03:20 V. Sulkosky BCM Log Parity run 12238
  • 289173 09/12/09 02:13 V. Sulkosky Compton HV Tripped
  • 289171 09/12/09 02:08 Xinhu Yan Target alarm
  • 289169 09/12/09 01:55 silwal spot++ on LH2, 6x4.5 mm
  • 289164 09/12/09 01:50 silwal spot++, holey target, right on the center, close enough..
  • 289163 09/12/09 01:46 V. Sulkosky Harp Scans after Compton Tune
  • 289160 09/12/09 01:32 silwal spot ++ with holey target
  • 289157 09/12/09 00:12 Xinhu Yan Target Screens
  • 289156 09/12/09 00:04 silwal swing shift summary
  • Friday

  • 289154 09/11/09 23:38 silwal hall A general tools, right arm Q2 still down
  • 289153 09/11/09 23:36 Kiadtisak Screenshot for Target Strip Charts
  • 289152 09/11/09 23:35 Kiadtisak Target Screen
  • 289151 09/11/09 23:30 dalton bpm12x saturation cut too high
  • 289150 09/11/09 22:59 O.Glamazdin Moller results
  • 289149 09/11/09 21:40 Jack (Techs) Right Q2 - latest trip
  • 289148 09/11/09 20:49 Kiadtisak Target moved to empty
  • 289147 09/11/09 19:57 silwal Q2 He level stipchart over the last couple of days, lots of spikes in flow recently
  • 289146 09/11/09 19:09 A. Camsonne Right Q2 quenched again
  • 289145 09/11/09 19:01 Kiadtisak Target moved to carbon hole targetr
  • 289144 09/11/09 18:23 A. Camsonne Compton computer network adapter dowm - computer rebooted
  • 289143 09/11/09 17:52 silwal Q2 reset, quenched again, He and N2 levels fine
  • 289142 09/11/09 17:45 Kiadtisak Target moved to empty
  • 289141 09/11/09 17:21 silwal mistake earlier, alexander had turned off the magnets. tried to bring them up, right arm Q2 quenched
  • 289140 09/11/09 17:04 silwal magnets down
  • 289139 09/11/09 16:22 Kiadtisak Target moved to optics
  • 289138 09/11/09 16:11 Kiadtisak Screenshot for Target Strip Charts
  • 289137 09/11/09 16:09 Kiadtisak Screenshots for target gui
  • 289136 09/11/09 16:09 jroche day shift summary
  • 289134 09/11/09 15:34 fansler/ewin Cryo/Power Supply Faults
  • 289121 09/11/09 13:42 M Friend Compton Asymmetry
  • 289118 09/11/09 13:34 jroche Heidi will access the Hall to look at LQ1
  • 289117 09/11/09 13:24 jroche beam tuning for at least one more hour
  • 289116 09/11/09 13:22 jroche Left Q1 He empty
  • 289115 09/11/09 13:20 A. Camsonne Compton computer crashed
  • 289114 09/11/09 13:04 Jeong Han Screenshot for Target Strip Charts
  • 289113 09/11/09 13:00 Jeong Han tgtgui screenshots
  • 289112 09/11/09 12:47 silwal modified .crl for LED setup in the left arm
  • 289107 09/11/09 12:10 R. Michaels write busy to ROC2 (HRS)
  • 289104 09/11/09 11:32 Nuruzzaman Beam Modulation turned off due to low current
  • 289097 09/11/09 10:03 dalton Redo test
  • 289096 09/11/09 09:57 jroche data taking while beam tune for Compton in progress
  • 289095 09/11/09 09:50 jroche 5 minutes at 50 uA to asses lumi asymmetry widths
  • 289094 09/11/09 09:39 R. Michaels re: raster X vs Y
  • 289091 09/11/09 09:33 dalton Rapid current change test
  • 289090 09/11/09 09:31 R. Michaels cleanup HapSimple
  • 289089 09/11/09 09:13 A. Camsonne Compton background rates - GSO in beam position
  • 289088 09/11/09 09:05 dalton Lumi width huge
  • 289086 09/11/09 08:25 V. Sulkosky Beam Profile Check with Harps
  • 289084 09/11/09 08:18 V. Sulkosky Explanation of ep Ion Chamber Trips
  • 289083 09/11/09 08:14 V. Sulkosky Re: Lumi Width not Asymmetry
  • 289082 09/11/09 08:06 paul Raster is back to 11x8.5
  • 289081 09/11/09 08:05 bogdanw/V. S Owl shift summary
  • 289079 09/11/09 07:55 V. Sulkosky Raster Signal Check L-HRS
  • 289078 09/11/09 07:53 dalton Target screenshots
  • 289077 09/11/09 07:52 dalton Target heater hitting floor hard
  • 289076 09/11/09 07:50 R. Michaels HapSimple and dat. (dot)
  • 289075 09/11/09 07:29 dalton Raster at 11 x 8.5
  • 289069 09/11/09 04:56 V. Sulkosky Beam Tuning Continues
  • 289068 09/11/09 04:40 dalton Optics target for tuning
  • 289067 09/11/09 04:22 Paul Adjusted the Raster Size - 11x7.5 (MCC) - 5.0x4.6 on target
  • 289062 09/11/09 03:33 dalton No data lost
  • 289061 09/11/09 03:33 dalton No data lost
  • 289060 09/11/09 03:17 silwal Parity DAQ hungup, has been resolved
  • 289059 09/11/09 03:17 silwal Parity DAQ hungup, has been resolved
  • 289051 09/11/09 01:59 bogdanw BPM, BCM
  • 289050 09/11/09 01:58 bogdanw Hall A tools
  • 289049 09/11/09 01:36 dalton Target Empty then back to H2
  • 289047 09/11/09 01:35 dalton spot++ this time correctly
  • 289045 09/11/09 01:02 dalton Target moved to H2
  • 289043 09/11/09 00:59 dalton Spot++
  • 289039 09/11/09 00:03 jpchen swing shift summary
  • Thursday

  • 289037 09/10/09 23:29 jpchen cryo event related to the left Q1 level drop
  • 289036 09/10/09 23:07 Navaphon cryotarget charts
  • 289035 09/10/09 23:06 Navaphon cryotarget
  • 289034 09/10/09 23:01 jpchen All magnets fine now
  • 289033 09/10/09 22:59 jpchen Hall in beam permit, beam tuning now
  • 289032 09/10/09 22:59 jpchen Right Q3 reset from the hall, ok
  • 289030 09/10/09 22:01 adaq cryoflow
  • 289029 09/10/09 22:00 jpchen right Q3 tripped. Q1
  • 289028 09/10/09 21:43 jpchen Q1_left tripped due to He level 10%
  • 289023 09/10/09 19:44 jpchen Hall access activities finished
  • 289004 09/10/09 18:12 Ramesh/Xiaoy access activity (pvdis work)
  • 289003 09/10/09 18:00 R. Subedi LHRS clock 1024 Hz
  • 288997 09/10/09 17:28 A. Camsonne Helicity back to delayed mode
  • 288995 09/10/09 16:56 A. Camsonne Helicity in time right now for tests
  • 288994 09/10/09 16:41 R. Michaels ADC18 DAC glitch cut required
  • 288980 09/10/09 16:00 Holmstrom Day SHift Summary
  • 288958 09/10/09 15:01 R. Michaels clock (T8) cleanup, R-HRS
  • 288952 09/10/09 14:13 Nuruzzaman Hall-A Tools screenshot
  • 288951 09/10/09 14:11 hyao target charts
  • 288950 09/10/09 14:10 hyao target tool screen
  • 288949 09/10/09 13:06 Ahmed/Mark Large difference of asymmetry
  • 288919 09/10/09 11:03 hyao target to empty
  • 288913 09/10/09 10:46 dalton Further monitor saturation complications
  • 288905 09/10/09 09:19 A. Camsonne Hall A in controlled access
  • 288904 09/10/09 08:07 bogdanw shift summary
  • 288902 09/10/09 08:00 silwal JT valve closed ~ 1.5 % to set the total A+C flow of 15K below 25 g/s
  • 288898 09/10/09 07:35 paul Target Controller Screenshot
  • 288897 09/10/09 07:33 bogdanw happex feedback monitor
  • 288892 09/10/09 05:35 R. Michaels may try to use vx5100 for roc1/2
  • 288885 09/10/09 04:19 silwal opened JT valve by 1 %
  • 288884 09/10/09 03:49 silwal cryo stripcharts
  • 288883 09/10/09 03:48 silwal cryo snapshot
  • 288879 09/10/09 02:36 bogdanw BPM, BCM strips
  • 288877 09/10/09 02:35 bogdanw Hall A tools
  • 288871 09/10/09 01:52 A. Camsonne Compton CODA buffer full
  • 288869 09/10/09 01:49 paul Run 12200
  • 288821 09/10/09 00:08 Kai HRS ps8 (clock) change to 20
  • 288819 09/10/09 00:05 hyao Swing shift summary
  • 288818 09/10/09 00:04 hyao ps8=20 for L,R arms
  • Wednesday

  • 288816 09/09/09 23:50 hyao beam is back.
  • 288795 09/09/09 21:48 hyao Controlled Access
  • 288793 09/09/09 21:26 R. Michaels ETROC firmware news
  • 288792 09/09/09 21:08 Navaphon target power and temperature
  • 288791 09/09/09 21:06 hyao Beam down
  • 288790 09/09/09 21:06 adaq Target screen shot
  • 288785 09/09/09 20:39 hyao BCM chart
  • 288784 09/09/09 20:39 hyao BPM chart
  • 288783 09/09/09 20:38 hyao hall a tool screen
  • 288782 09/09/09 20:02 hyao beam back. Taking data
  • 288780 09/09/09 19:37 hyao Beam down.
  • 288775 09/09/09 18:28 O.Glamazdin HAPPEX CODA crash
  • 288773 09/09/09 18:17 hyao beam is back
  • 288771 09/09/09 17:51 hyao MCC turn off the beam.
  • 288767 09/09/09 16:50 A. Camsonne Compton FADC crate power cycled
  • 288764 09/09/09 16:01 Saha Shift Summay
  • 288762 09/09/09 15:48 mercado Target Screen Shots
  • 288761 09/09/09 15:44 mercado Target Stripcharts
  • 288758 09/09/09 15:23 Saha Hall A Tools
  • 288757 09/09/09 15:22 Saha BPM Striptool
  • 288756 09/09/09 15:21 saha BCM Striptool
  • 288751 09/09/09 13:20 rsholmes Compton gains (formatted)
  • 288750 09/09/09 13:20 rsholmes Compton gains
  • 288748 09/09/09 12:49 apar Halfwave plate OUT, Slug 5
  • 288746 09/09/09 12:39 A. Camsonne Compton Electron detector in beam position
  • 288743 09/09/09 12:19 rsholmes Compton spectra from counting
  • 288742 09/09/09 11:49 M Friend Compton Base Current Instabilities
  • 288732 09/09/09 08:30 mercado Compton HV trip.
  • 288730 09/09/09 08:06 bogdanw shift summary
  • 288727 09/09/09 07:09 bogdanw BPM, BCM
  • 288726 09/09/09 07:08 bogdanw Hall A tools
  • 288723 09/09/09 06:07 Abdurahim Target strip charts
  • 288722 09/09/09 06:06 Abdurahim Target screen shots
  • 288717 09/09/09 05:03 A. Camsonne Compton cavity off for background check
  • 288716 09/09/09 04:53 A. Camsonne Compton DAQ flaky at high rate
  • 288713 09/09/09 04:09 Abdurahim Target strip charts
  • 288712 09/09/09 04:08 Abdurahim Target screen shots
  • 288707 09/09/09 02:06 Abdurahim Target strip charts
  • 288706 09/09/09 02:06 Abdurahim Target screen shots
  • 288703 09/09/09 01:09 paul General Tools Screenshot
  • 288702 09/09/09 01:08 paul BPM screenshot
  • 288701 09/09/09 01:07 paul BCM screeshot
  • 288700 09/09/09 00:49 paul After gain adjustment the BPM12 reading went up to 140000
  • 288696 09/09/09 00:44 paul We stopped the run #12163 for gain adjustment on BPM12.
  • 288692 09/09/09 00:00 X. Zhan Shift Summary
  • Tuesday

  • 288689 09/08/09 22:52 rsholmes 18 bit BPM12: there are always complications
  • 288688 09/08/09 22:43 mercado More about bmp12 change.
  • 288687 09/08/09 22:42 apar Test of flipper halog entry: Halfwave plate IN, Slug 4
  • 288686 09/08/09 22:40 mercado HAPPEX DAQ: Changed bmp12 to 18-bit ADC
  • 288680 09/08/09 21:25 R. Michaels 18-bit ADC and BPM10.
  • 288677 09/08/09 20:54 A. Camsonne Compton events lost
  • 288676 09/08/09 20:32 Xiaohui Hall A Tool
  • 288673 09/08/09 20:08 M Friend Compton Spectra
  • 288672 09/08/09 20:01 O.Glamazdin spot++
  • 288671 09/08/09 19:34 Chiranjib Target charts III
  • 288670 09/08/09 19:33 Chiranjib Target charts II
  • 288669 09/08/09 19:33 Chiranjib Target charts I
  • 288665 09/08/09 19:07 mercado HAPPEX DAQ: Added another 18-bit Voltage ADC to the Counting House
  • 288663 09/08/09 18:41 Mark/Ahmed Large position difference. Run 11996
  • 288661 09/08/09 17:47 Xiaohui Beam is away
  • 288660 09/08/09 17:27 Xiaohui Target--->LH loop 3
  • 288654 09/08/09 16:39 Chiranjib Target ---> Empty
  • 288651 09/08/09 16:11 dalton Charge feedback history
  • 288650 09/08/09 16:03 Saha Shift Summary
  • 288648 09/08/09 15:34 hyao target charts
  • 288647 09/08/09 15:34 hyao target tool screen
  • 288646 09/08/09 15:30 buddhini beam modulation
  • 288645 09/08/09 15:27 A. Camsonne Compton rates very instable
  • 288641 09/08/09 14:55 buddhini beam modulation
  • 288640 09/08/09 14:18 A. Camsonne Restored U1 cabling
  • 288636 09/08/09 13:33 R. Subedi Access acitvity (pvdis only)
  • 288623 09/08/09 13:00 A. Camsonne Moved Compton IOC on the floor
  • 288620 09/08/09 12:51 A. Camsonne Stole a PS758 from BigBite cabinet
  • 288607 09/08/09 12:03 rsholmes Compton photon counting polarizations
  • 288605 09/08/09 12:00 A. Camsonne Compton scaler BCM check
  • 288591 09/08/09 11:37 R. Michaels L-HRS DAQ restored
  • 288577 09/08/09 11:04 A. Camsonne Hall in controlled access until 3PM Full survey complete
  • 288574 09/08/09 11:00 A. Camsonne Compton V260 scaler reset signal coming from NIM/ECL FAN converter stuck at 1
  • 288547 09/08/09 08:09 bogdanw shift summary
  • 288545 09/08/09 08:03 Abdurahim Compton cavity is Off
  • 288544 09/08/09 08:01 Abdurahim Target screen shots
  • 288543 09/08/09 08:01 Abdurahim Target strip charts
  • 288542 09/08/09 07:51 paul Dropped current to 10uA for photon line tune
  • 288537 09/08/09 07:15 bogdanw Compton HV
  • 288536 09/08/09 06:50 Abdurahim Compton DAQ problem
  • 288533 09/08/09 06:24 bogdanw Compton HV, need expert
  • 288532 09/08/09 06:23 Abdurahim Compton machine is rebooted
  • 288531 09/08/09 06:17 Paul BPM sceenshot
  • 288530 09/08/09 06:16 Paul BCMs stripcharts
  • 288529 09/08/09 06:15 Paul Hall A Genral Tools Screenhshot
  • 288528 09/08/09 06:04 Abdurahim Target strip charts
  • 288527 09/08/09 06:04 Abdurahim Target screen shots
  • 288526 09/08/09 05:56 Paul Recent CODA errors ... Switched to "HAPEX" configuration
  • 288514 09/08/09 04:42 A. Camsonne Standard DAQ issue
  • 288511 09/08/09 04:32 A. Camsonne GSO HV trip
  • 288508 09/08/09 04:06 Abdurahim Target strip charts
  • 288507 09/08/09 04:05 Abdurahim Target screen shot
  • 288506 09/08/09 03:38 A. Camsonne GSO HV trip
  • 288502 09/08/09 03:02 Abdurahim Compton DAQ problem
  • 288501 09/08/09 02:49 Abdurahim standard Compton DAQ problem
  • 288498 09/08/09 02:04 bogdanw beam modulation
  • 288497 09/08/09 02:03 Abdurahim Target screen shots
  • 288496 09/08/09 02:03 Abdurahim Target strip charts
  • 288495 09/08/09 01:48 bogdanw beam position, current strips
  • 288494 09/08/09 01:46 bogdanw Hall A tools, Feedback
  • Monday

  • 288488 09/07/09 23:55 Xiaohui Zhan Shift Summary
  • 288487 09/07/09 23:48 Xiaohui Hall A Tool
  • 288484 09/07/09 23:18 A. Camsonne Compton electron detector out
  • 288483 09/07/09 23:16 A. Camsonne Run was 19015
  • 288482 09/07/09 23:10 A. Camsonne Compton Electron detector spectrum at 0 mm
  • 288481 09/07/09 23:06 A. Camsonne Compton DAQ switched to photon only for production
  • 288477 09/07/09 22:01 A. Camsonne ETROC latest standard version reloaded
  • 288476 09/07/09 21:50 A. Camsonne ETROC photon trigger only firmware loaded
  • 288475 09/07/09 21:39 Kalyan loop 1 read back problem ?
  • 288473 09/07/09 21:35 Xiaohui Beam current chart
  • 288471 09/07/09 20:39 A. Camsonne Compton electron detector in beam position
  • 288470 09/07/09 20:35 A. Camsonne Compton electron detector currents out of the beam
  • 288466 09/07/09 20:19 Kalyan Opened coolant JT valve by 1%
  • 288465 09/07/09 20:12 Kalyan Target strip charts
  • 288464 09/07/09 20:12 Kalyan Target screen
  • 288461 09/07/09 19:05 mercado HAPPEX DAQ - Added pedestal calibration to boot files
  • 288460 09/07/09 18:50 Chiranjib BPM12 saturation
  • 288458 09/07/09 18:37 Xiaohui Beam is back
  • 288457 09/07/09 18:28 Xiaohui Beam is away
  • 288455 09/07/09 18:14 Xiaohui Beam current
  • 288453 09/07/09 17:37 R. Michaels run 24668 no FB gates
  • 288450 09/07/09 17:21 Xiaohui MCC is ready to send the beam back
  • 288449 09/07/09 17:13 R. Michaels re : LHRS DAQ
  • 288448 09/07/09 17:07 Xiaohui Beam Off
  • 288446 09/07/09 16:58 Xiaohui Started production
  • 288445 09/07/09 16:55 Xiaohui Hall A Tool
  • 288442 09/07/09 16:41 Quinn Changes in CMU compton DAQ
  • 288440 09/07/09 16:37 A. Camsonne ETROC latest firmware version loaded
  • 288437 09/07/09 16:17 mercado Beam Energy Monitor
  • 288436 09/07/09 16:16 mercado, Kai Day Shift Summary
  • 288433 09/07/09 16:11 R. Subedi Left HRS problem NOT with L1A but with spot++
  • 288430 09/07/09 16:01 buddhini run summary
  • 288429 09/07/09 16:00 A. Camsonne Compton ETROC FPGA programmer on hareboot4 port 5
  • 288428 09/07/09 15:32 mercado Striptools
  • 288427 09/07/09 15:31 mercado Hall A Tools
  • 288426 09/07/09 15:19 mercado Beam Down
  • 288423 09/07/09 15:10 hyao target tool screen
  • 288422 09/07/09 15:09 hyao target charts
  • 288421 09/07/09 15:06 M Friend Compton FADC DAQ Asymmetries
  • 288418 09/07/09 14:04 hyao average haH2_P_Ballast_R drops
  • 288415 09/07/09 12:30 hyao spike of H2_P_Ballast_R
  • 288412 09/07/09 12:20 hyao haH2_P_Ballast_R yellow warning
  • 288411 09/07/09 12:09 hyao Open JT valve 1%
  • 288410 09/07/09 11:55 M Friend Beam Back in Position
  • 288409 09/07/09 11:40 M Friend Compoton Beam Position Problem
  • 288408 09/07/09 11:27 Kai BPM12 saturates again in #12079. Luis is back, and on shift now.
  • 288401 09/07/09 10:02 Kai online replay of new run #12078 looks good
  • 288400 09/07/09 09:53 Kai compton #18989 ended failed, reset and start a new run # 18990
  • 288399 09/07/09 09:51 buddhini bpm saturation
  • 288398 09/07/09 09:48 rsholmes BPM12 again
  • 288397 09/07/09 09:37 R. Michaels re: fixing L-HRS
  • 288396 09/07/09 09:32 Kai But Online Reply shows the BMP12 totally saturated in Run # 12076.
  • 288395 09/07/09 09:32 adaq bpm saturate
  • 288392 09/07/09 09:09 Kai Good Run # 12076, ended normally
  • 288389 09/07/09 08:32 Kai Run 12075 end failed.
  • 288387 09/07/09 08:07 bogdanw shift summary
  • 288383 09/07/09 07:20 paul hall A general tools screen
  • 288382 09/07/09 07:19 paul bpm screen
  • 288381 09/07/09 07:19 paul bcm screens
  • 288378 09/07/09 06:56 dalton Target stripcharts
  • 288375 09/07/09 05:13 dalton compton coda reboot
  • 288374 09/07/09 05:13 paschke L-HRS spectrometer daq down due to cable snagged
  • 288371 09/07/09 05:09 paschke BPM saturation cut applied to default .db's
  • 288370 09/07/09 05:08 Paul Beam energy monitor screenshot
  • 288369 09/07/09 04:27 Paul General Tools Screenshot
  • 288368 09/07/09 04:22 paul Control Screenshot
  • 288361 09/07/09 02:12 paul Feedback monitor screenshot
  • 288360 09/07/09 01:50 A. Camsonne Compton BCM puzzle
  • 288359 09/07/09 01:49 dalton Hydrogen pressure alarms again
  • 288356 09/07/09 01:11 A. Camsonne Compton electron detector current in garage position
  • 288355 09/07/09 01:01 paul BCM stripchart Screenshot
  • 288354 09/07/09 01:00 paul BPMs stripchart Screenshot
  • 288353 09/07/09 00:59 paul Haa A general Tools Screenshot
  • 288350 09/07/09 00:05 silwal swing shift summary
  • 288349 09/07/09 00:00 silwal beam energy monitor
  • Sunday

  • 288348 09/06/09 23:39 Abdurahim Hydrogen pressure is stable now
  • 288346 09/06/09 23:29 silwal beam cur/position stipcharts
  • 288344 09/06/09 23:28 silwal hall A tools
  • 288343 09/06/09 23:23 Abdurahim Target screen shots
  • 288342 09/06/09 23:22 Abdurahim The hydrogen pressure is keep dropping
  • 288341 09/06/09 23:16 silwal pedestal corrected det1/det2 rms
  • 288340 09/06/09 23:15 Abdurahim strip charts correlated with target alarm
  • 288339 09/06/09 23:10 Abdurahiim target alarm
  • 288338 09/06/09 22:39 silwal compton HV trip
  • 288335 09/06/09 22:16 Abdurahim Target strip charts
  • 288334 09/06/09 22:16 Abdurahim Target screen shots
  • 288333 09/06/09 22:08 silwal beam cur/position, dithering on
  • 288332 09/06/09 21:44 A. Camsonne Rebooted Compton Computer
  • 288327 09/06/09 20:39 paschke FastFeedback and BPM12 problems tonight
  • 288324 09/06/09 20:27 paschke dithering on
  • 288323 09/06/09 20:13 paschke Dithing looks good, to be turned on
  • 288322 09/06/09 20:01 silwal beam cur/ position stripcharts
  • 288321 09/06/09 19:58 Abdurahim Target strip charts
  • 288320 09/06/09 19:58 A. Camsonne EP trips
  • 288319 09/06/09 19:57 Abdurahim Target screen shots
  • 288318 09/06/09 19:38 A. Camsonne Radiation spikes
  • 288317 09/06/09 19:36 silwal System FFB on, bpm12 gain fine
  • 288314 09/06/09 19:30 A. Camsonne Power cycled Compton DAQ and ETROC
  • 288313 09/06/09 19:28 A. Camsonne Compton rates ok
  • 288312 09/06/09 19:12 silwal FFB snapshot after bcm12 gains fixed
  • 288311 09/06/09 18:41 silwal beam energy monitor
  • 288309 09/06/09 18:29 silwal left arm det HV changed to -1590
  • 288303 09/06/09 18:04 silwal hall A tools snapshot
  • 288300 09/06/09 17:58 Abdurahim Target strip charts
  • 288299 09/06/09 17:57 Abdurahim Target screen shots
  • 288297 09/06/09 17:46 R. Michaels adc18_stop function added to lib
  • 288295 09/06/09 17:29 silwal added adc18_loaddac & adc18_setsample to boot scripts
  • 288294 09/06/09 17:18 A. Camsonne Compton reduced electron detector buffer size
  • 288293 09/06/09 17:14 A. Camsonne Electron detector in garage position
  • 288291 09/06/09 16:42 Kalyan Beam is back, spot++
  • 288290 09/06/09 16:41 silwal turned the right arm S0 on/off
  • 288283 09/06/09 16:12 A. Camsonne Runs for comparison
  • 288282 09/06/09 15:59 Z. Ahmed Day shift summary
  • 288279 09/06/09 15:03 R. Michaels Blaster network tests
  • 288278 09/06/09 14:47 R. Michaels advice re: Inj Crate
  • 288277 09/06/09 14:46 Z. Ahmed Half-wave plate is out.
  • 288276 09/06/09 14:44 Z. Ahmed Half-wave plate is out.
  • 288275 09/06/09 14:21 R. Michaels L-HRS spectr. DAQ no gates
  • 288274 09/06/09 13:01 A. Camsonne Compton BCM
  • 288273 09/06/09 12:21 M Friend Compton HV Trip
  • 288272 09/06/09 12:19 Z. Ahmed Strip tool
  • 288271 09/06/09 12:18 Z. Ahmed. Hall A General Tools.
  • 288270 09/06/09 12:17 hyao target chart
  • 288269 09/06/09 12:16 hyao target tools screen
  • 288268 09/06/09 11:38 Z. Ahmed Accelerator is down.
  • 288267 09/06/09 11:34 rsholmes Compton BCM (attempt formatting)
  • 288266 09/06/09 11:33 rsholmes Compton BCM
  • 288265 09/06/09 10:39 Z. Ahmed. Accelerator is Down.
  • 288263 09/06/09 10:23 Z. Ahmed Target heater.
  • 288258 09/06/09 08:44 dalton Slug summaries
  • 288256 09/06/09 08:01 dalton Runbird screenshot
  • 288253 09/06/09 07:39 dalton FFB back on
  • 288250 09/06/09 06:54 dalton Changes to database
  • 288243 09/06/09 04:29 Piotr Target charts
  • 288242 09/06/09 04:27 Piotr Target screens
  • 288241 09/06/09 04:06 silwal injector crate hungup
  • 288240 09/06/09 03:58 paschke Compton HV GSO trip
  • 288239 09/06/09 03:56 silwal adc18 gains preset in the boot files for counting house, left and right arm crates
  • 288238 09/06/09 03:56 silwal adc18 gains preset in the boot files for counting house, left and right arm crates
  • 288237 09/06/09 03:54 silwal adc18 gains preset in the boot files for counting house, left and right arm crates
  • 288234 09/06/09 03:38 paschke DAQ trouble and L-HRS parity daq crate reset
  • 288229 09/06/09 01:44 dalton Raster still good
  • 288227 09/06/09 01:41 dalton 100 uA without FFB
  • 288213 09/06/09 00:43 M Friend Compton HV Trip
  • 288212 09/06/09 00:42 dalton spot_R, 100 uA, PS8=1
  • 288208 09/06/09 00:29 silwal compton HV trip
  • 288205 09/06/09 00:21 silwal hall A general tools
  • 288204 09/06/09 00:14 M Friend Moved Compton FADC DAQ BCM
  • 288202 09/06/09 00:03 silwal swing shift summary
  • Saturday

  • 288201 09/05/09 23:57 Abdurahim Target strip charts at the end of swing shift
  • 288200 09/05/09 23:56 Abdurahim Target screen shot at the end of swing shift
  • 288196 09/05/09 23:46 paschke spot++ broken
  • 288189 09/05/09 23:25 silwal Counting house adc gain @ 3,0-should have been 2,0
  • 288187 09/05/09 23:22 paschke adc gains wrong AGAIN
  • 288186 09/05/09 23:20 silwal HallA FFB RF turn on problem
  • 288183 09/05/09 22:44 silwal alarm on due to VDC, despite the VDC's being off
  • 288182 09/05/09 22:18 Abdurahim Traget strip charts
  • 288181 09/05/09 22:17 Abdurahim Target screen shots
  • 288177 09/05/09 21:59 O.Glamazdin Moller results
  • 288176 09/05/09 21:44 Abdurahim Target Status
  • 288175 09/05/09 20:45 paschke repeating plot of dithering cycle
  • 288174 09/05/09 20:40 paschke Surprising! results of dithering system test
  • 288173 09/05/09 20:36 Abdurahim Target strip charts
  • 288172 09/05/09 20:35 Abdurahim Target screen shots
  • 288171 09/05/09 20:07 paschke Dither test concluded
  • 288170 09/05/09 19:20 silwal beam enerty monitor snapshot
  • 288169 09/05/09 19:09 Abdurahim Target Status
  • 288162 09/05/09 18:00 Abdurahim Target strip charts
  • 288161 09/05/09 17:58 Abdurahim Target screen shots
  • 288159 09/05/09 17:24 silwal no beam since 5:10 pm
  • 288153 09/05/09 16:02 mercado Day Shift Summary
  • 288152 09/05/09 15:41 mercado Beam Monitor
  • 288151 09/05/09 15:24 mercado Beam Energy Monitor
  • 288148 09/05/09 15:04 mercado Hall A Tools
  • 288147 09/05/09 15:03 M Friend Moved Compton Collimator
  • 288146 09/05/09 14:51 A. Camsonne Compton ROC1 error message at boot
  • 288145 09/05/09 14:35 mercado new DAC sync between Counting House and Right Arm
  • 288143 09/05/09 14:23 paschke Summary of controlled access activity - DAQ timing checked
  • 288142 09/05/09 14:22 mercado Controlled Access over
  • 288141 09/05/09 14:18 A. Camsonne Amplifier discriminator board in for plan A and Compton electron detector in beam position
  • 288139 09/05/09 13:53 A. Camsonne Frequent reason for Compton ROC crashes
  • 288137 09/05/09 13:15 hyao target charts
  • 288136 09/05/09 13:13 hyao target screens
  • 288135 09/05/09 12:45 mercado HAPPEX DAQ Configuration Change
  • 288127 09/05/09 11:18 mercado Half-wave plate IN
  • 288126 09/05/09 11:07 silwal pvdis_oversample1.ped and pvdis.dacnoise created to keep createDB from reading default dacnoise and ped
  • 288125 09/05/09 10:32 paschke changed analysis default helicity map
  • 288122 09/05/09 09:44 paschke PAN .db template trimmed
  • 288119 09/05/09 09:41 hyao Loop3 Heater Warning
  • 288118 09/05/09 09:35 R Michaels re: integ. time diff ?
  • 288115 09/05/09 08:31 A. Camsonne Analog signal plane A self triggered
  • 288114 09/05/09 08:15 A. Camsonne High luminosity Compton ? ( BCM looks weird )
  • 288113 09/05/09 08:02 ZAFAR AHMED2 OWL Shif Summary
  • 288110 09/05/09 07:05 A. Camsonne Compton EPICS reboot
  • 288109 09/05/09 06:58 Piotr JT valve opened 1%
  • 288106 09/05/09 06:42 A. Camsonne Compton electron detector analog signal
  • 288105 09/05/09 06:31 Piotr Target screens
  • 288104 09/05/09 06:30 Piotr Target charts
  • 288103 09/05/09 06:02 Z. Ahmed Integration gate time same.
  • 288102 09/05/09 06:00 A. Camsonne Compton Electron detector current at position 2mm
  • 288101 09/05/09 05:58 A. Camsonne Compton coincidence electron data for run 18907
  • 288100 09/05/09 05:53 Z. Ahmed Integration gate time changed
  • 288099 09/05/09 05:52 Higinbotham Shifts of Target Ladder
  • 288096 09/05/09 05:51 A. Camsonne Moving electron detector to position 2mm
  • 288093 09/05/09 04:36 adaq run 12013
  • 288089 09/05/09 03:43 R. Michaels auto abu fixed (again)
  • 288086 09/05/09 02:09 Z. Ahmed HV left and right
  • 288085 09/05/09 02:09 Piotr heater power fluctuates
  • 288083 09/05/09 02:08 Z. Ahmed Hall A General Tools
  • 288081 09/05/09 01:34 D.Wang Graphs
  • 288080 09/05/09 01:19 D.Wang compton efficiency
  • 288075 09/05/09 00:01 Fatiha Shift Summary
  • 288073 09/05/09 00:00 aniol target screen
  • Friday

  • 288072 09/04/09 23:56 aniol target charts
  • 288069 09/04/09 23:09 Fatiha Keep an eye on the ABU
  • 288063 09/04/09 20:30 Fatiha More stable beam currents
  • 288062 09/04/09 20:25 dalton Put a run summery in halog?
  • 288061 09/04/09 20:21 dalton 50% pair efficiency problem
  • 288057 09/04/09 19:59 Fatiha Happex Feedback Monitor
  • 288054 09/04/09 19:14 M Friend Compton Photon Beam Position
  • 288053 09/04/09 19:06 apar Test of prompt summary halog: Number 2
  • 288052 09/04/09 19:05 R. Michaels auto abu enabled at 5:30 pm
  • 288051 09/04/09 19:03 apar Test of prompt summary halog
  • 288050 09/04/09 18:31 Fatiha Beam current fluctuations
  • 288047 09/04/09 18:11 R. Michaels requested spot++ check run (if RC agrees)
  • 288046 09/04/09 17:35 Fatiha Spot++ on LH2 target @ 100 uA
  • 288040 09/04/09 17:13 R. Michaels spot++
  • 288038 09/04/09 17:01 R. Michaels Runs with regular Alum dummy
  • 288035 09/04/09 16:56 R. Michaels Alum. dummy hole tgt part II
  • 288034 09/04/09 16:48 Fatiha Beam current screen
  • 288027 09/04/09 16:29 R. Michaels Dummy hole alum. tgt. part I
  • 288022 09/04/09 16:23 Fatiha Hall A Tool screen
  • 288019 09/04/09 16:16 R. Michaels Well centered hole now.
  • 288013 09/04/09 16:06 R. Michaels Hole after movement #1.
  • 288012 09/04/09 16:02 hyao Day Shift Summary
  • 288011 09/04/09 16:01 R. Michaels High statistics hole pic
  • 288006 09/04/09 15:53 R. Michaels C12 Holey target.
  • 287999 09/04/09 15:43 R. Michaels Beam current scan
  • 287998 09/04/09 15:41 W. Deconinck Carbon hole target
  • 287989 09/04/09 15:15 mcnulty alignment and detector ADC attenuation from runs 24650/4459
  • 287988 09/04/09 15:12 R. Michaels Rate problem understood
  • 287987 09/04/09 15:07 R. Michaels Qsq runs
  • 287980 09/04/09 15:00 R. Michaels Rate discrepancy
  • 287977 09/04/09 14:55 M Friend Compton FADC DAQ Plots
  • 287976 09/04/09 14:52 A. Camsonne High statistics Compton electron coincidence run
  • 287975 09/04/09 14:49 mcnulty Going to sampling mode for Q2 meas.
  • 287974 09/04/09 14:47 R. Michaels funny looking spot++ for LH2
  • 287973 09/04/09 14:45 W. Deconinck Measurement programme: Q2 measurement, hole search
  • 287970 09/04/09 14:42 Hall A worke ROC31 disconnected
  • 287968 09/04/09 14:30 R. Michaels DAQ switchover checklist
  • 287967 09/04/09 14:27 mercado Noticed weird pedestal jump/drop in 18-bit Qweak ADC (bcm7)
  • 287966 09/04/09 14:26 hyao Hall A Tools
  • 287965 09/04/09 14:25 W. Deconinck Target screen capture
  • 287964 09/04/09 14:22 mercado Even Better Double Difference
  • 287963 09/04/09 13:49 R. Michaels halog problem / solution
  • 287960 09/04/09 13:37 M Friend Compton FADC DAQ Update
  • 287959 09/04/09 13:33 A. Camsonne Asked MCC to check Compton Background
  • 287953 09/04/09 13:25 A. Camsonne First Compton photon electron coincidence
  • 287952 09/04/09 13:24 W. Deconinck Target status: all is well.
  • 287951 09/04/09 13:02 A. Camsonne Edet in position at 4mm
  • 287950 09/04/09 12:50 hyao Beam Off for 20 minutes
  • 287947 09/04/09 12:19 mercado BCM double difference
  • 287946 09/04/09 11:56 mercado Double difference of BCM1 and BCM6
  • 287945 09/04/09 11:37 mercado new BCM6 = BCM2 + 18-bit ADC
  • 287944 09/04/09 11:31 R. Michaels Central beam position ?
  • 287943 09/04/09 11:31 M Friend Compton FADC DAQ Progress
  • 287940 09/04/09 11:20 R. Michaels FADCScaler to top of list
  • 287939 09/04/09 11:04 M Friend Correct Compton CODA Configuration
  • 287938 09/04/09 10:59 R. Michaels Cavities work now !
  • 287935 09/04/09 10:02 hyao Compton Run Start.
  • 287930 09/04/09 08:29 Kai Pan owl shift summary
  • 287928 09/04/09 08:24 Kai Kent called in, and asked to change Right arm Happex HV from -1700 to -1200
  • 287925 09/04/09 08:06 Kai 7:50am MCC called that they cannot reach 100uA, only 90uA is available for now.
  • 287924 09/04/09 08:05 Kai 7:45AM start production run #11991, though it is only 80uA
  • 287923 09/04/09 07:58 hoyoung Run list summary for owl shift
  • 287922 09/04/09 07:57 hoyoung 90 uA current
  • 287921 09/04/09 07:56 Kai Holey Target run and measurement of Happex scaler rate
  • 287918 09/04/09 07:47 Kalyan H2 temperature and pressure strip charts
  • 287917 09/04/09 07:45 Kalyan Target screens
  • 287899 09/04/09 06:52 Kai use new raster size 11*8.5 to do Holey target run and Happex counting run again
  • 287898 09/04/09 06:48 Kai Bob called in to explain spot++ problem. Problem solved!
  • 287897 09/04/09 06:45 Kalyan Special runs in counting mode
  • 287888 09/04/09 06:33 Kalyan Target was liquid hydrogen for spot size run, 4452
  • 287887 09/04/09 06:32 R. Michaels clarfications about spot++
  • 287886 09/04/09 06:30 Kai 5x4.5mm spot size
  • 287884 09/04/09 06:26 Kai read out of Happex scaler counting
  • 287874 09/04/09 06:09 Kai read out of Happex scaler counting
  • 287869 09/04/09 05:54 Kalyan Target to 25cm Liquid H2
  • 287868 09/04/09 05:53 Kai Turn VDC Off, Target move to LH2, prepare to measure rate in Happex-D.
  • 287861 09/04/09 05:42 Kai both arm are in counting mode, and HV of s0 are turned on in both arms
  • 287858 09/04/09 05:35 Kai Holey Target. Turn VDC on
  • 287857 09/04/09 05:33 Kai spot++ #4446, changed back to 11*8.5 raster size again, result still bad, decide to run in 7.7*8 at this moment
  • 287856 09/04/09 05:27 Kai spot++ for 7.7*8.0 is better, but not optimal
  • 287855 09/04/09 05:26 Kalyan Target moved to carbon hole
  • 287852 09/04/09 05:17 Kai spot++ result attached
  • 287851 09/04/09 05:15 Kai spot++ #4444. Result not good.
  • 287840 09/04/09 04:49 Kai Harp Scan result.
  • 287833 09/04/09 04:44 Kai MCC finished Harp Scan. Result good.
  • 287830 09/04/09 04:36 A. Camsonne High background rate for Compton
  • 287829 09/04/09 04:33 Kalyan Target moved to optics(carbon) target
  • 287828 09/04/09 04:30 Kai 4:05AM Asked MCC to do Harp Scan. Turn off the raster.
  • 287827 09/04/09 04:27 Kai 3:55AM Stable CW 10uA beam sent to the Hall
  • 287826 09/04/09 04:26 Kai 3:45AM status changed to beam permit.
  • 287825 09/04/09 04:25 Kai 3:20AM MPS inversion puzzle solved.
  • 287808 09/04/09 01:31 A. Camsonne Compton scope on halla-7
  • 287805 09/04/09 01:27 Kai still controlled access
  • 287802 09/04/09 01:06 A. Camsonne Modfication ETROC
  • 287796 09/04/09 00:16 A. Camsonne ETROC photon trigger switch with external trigger on input 6 now
  • 287794 09/04/09 00:07 Kai Owl shift start, hall in controlled access
  • Thursday

  • 287789 09/03/09 23:58 Fatiha Hall in controlled access
  • 287786 09/03/09 23:46 Aniol target screen
  • 287785 09/03/09 23:46 A. Camsonne Electron detector status
  • 287784 09/03/09 23:40 A. Camsonne Edet Currents no beam
  • 287783 09/03/09 23:32 A. Camsonne Compton electron detector activity summary
  • 287782 09/03/09 23:31 Fatiha Beam will be back around midnight
  • 287781 09/03/09 23:27 A, Camsonne Compton electron detector threshold scan
  • 287775 09/03/09 22:14 Fatiha Hall going to controled access
  • 287773 09/03/09 21:50 Fatiha Magnet screen
  • 287772 09/03/09 21:38 Fatiha Beam Status
  • 287771 09/03/09 21:37 M Friend Compton FADC DAQ
  • 287770 09/03/09 21:36 R. Michaels Possible MPS inversion (problem ?)
  • 287765 09/03/09 21:22 R. Michaels DAQ problem (fixed)
  • 287759 09/03/09 21:16 A. Camsonne Electron detector noise after grounding of board to box
  • 287754 09/03/09 20:18 Jack (Techs) Left Q3 trip function checked
  • 287753 09/03/09 20:14 Jack (Techs) Right Dipole - jumpers replaced and meter checked
  • 287751 09/03/09 20:06 Techs Checklist complete
  • 287733 09/03/09 18:54 R. Michaels New DAQ config: HapSimple
  • 287729 09/03/09 18:47 adaq New Datamap for cavity 3x(Dwon stream)
  • 287728 09/03/09 18:34 R. Subedi Remote reset of PVDIS crates
  • 287726 09/03/09 18:13 R. Michaels Problem: gains after reboot
  • 287723 09/03/09 18:08 A. Camsonne Compton electron detector pick-up noise
  • 287719 09/03/09 18:04 R. Michaels DAQ tests and status
  • 287689 09/03/09 16:33 rsholmes Wein angle
  • 287686 09/03/09 16:18 R. Michaels raster current readout restored
  • 287678 09/03/09 15:52 Higinbotham Shift Summary
  • 287675 09/03/09 15:42 Higinbotham Combined Energy and Phase Results
  • 287673 09/03/09 14:49 W. Deconinck Starting ramp-up of L Q2,3 (cycle) and L,R D
  • 287669 09/03/09 14:17 R. Michaels re: evtypebits
  • 287668 09/03/09 14:00 A. Camsonne Standard DAQ fixed followup
  • 287667 09/03/09 13:53 Z. Ahmed. Cavity 3x channel changed.
  • 287666 09/03/09 13:24 mcnulty Larm HapDet TDC spectrum looks good
  • 287665 09/03/09 13:13 mcnulty Larm s0 Larm s0 TDCs are OK for run 24599
  • 287662 09/03/09 12:44 W. Deconinck Target to empty position
  • 287654 09/03/09 12:04 mcnulty s0 ADC and TDC spectra for Right and Left Arms
  • 287651 09/03/09 11:57 W. Deconinck Kai Pan: day, 3rd: 5123 (pager)
  • 287650 09/03/09 11:53 W. Deconinck Restricted access -> TO required, shift leader and 3rd person on call
  • 287649 09/03/09 11:46 silwal can't telnet into hatsv18
  • 287647 09/03/09 11:13 silwal ROC31 hangup
  • 287642 09/03/09 08:11 mercado Owl Shift Summary
  • 287639 09/03/09 07:43 Kalyan another test of L-HRs happex detector
  • 287629 09/03/09 05:56 mercado BPM Striptool
  • 287628 09/03/09 05:55 mercado Hall A Tools
  • 287606 09/03/09 02:58 Kalyan L3 temperature and power strip charts
  • 287605 09/03/09 02:58 Kalyan Target Screen
  • 287604 09/03/09 02:35 Kalyan Run 24635 is for testing PMT base with HV -2100V
  • 287601 09/03/09 02:31 Kalyan Re: det1 (LHRS) with base H3/Bugs Bunny -- No ADC signal at -2100V
  • 287593 09/03/09 01:14 silwal PITA scan with IHWP out
  • 287589 09/03/09 00:20 M Friend Standard Compton DAQ fixed
  • 287587 09/03/09 00:18 mercado Beam is back
  • 287585 09/03/09 00:01 Fatiha Shift Summary
  • Wednesday

  • 287583 09/02/09 23:56 Fatiha Beam Down
  • 287581 09/02/09 23:50 Fatiha HWP is OUT (Happex Feedback Monitor)
  • 287580 09/02/09 23:48 Fatiha Half Wave Plate change from IN to OUT
  • 287578 09/02/09 23:32 dalton Dignity restored
  • 287577 09/02/09 23:31 Aniol JT valve - small increase
  • 287576 09/02/09 23:28 Fatiha DAQ is back
  • 287574 09/02/09 23:07 Aniol target screen
  • 287573 09/02/09 22:52 Fatiha DAQ issues
  • 287572 09/02/09 22:42 Fatiha Beam is back
  • 287571 09/02/09 21:59 dalton Laser signals without their pants on
  • 287570 09/02/09 21:53 silwal replay mismatch betn pan and pan_regress
  • 287569 09/02/09 21:51 Fatiha Beam status Update
  • 287567 09/02/09 21:21 adaq abnormal replay again?
  • 287566 09/02/09 21:13 Fatiha Beam down again
  • 287565 09/02/09 21:11 Fatiha Q3 back
  • 287563 09/02/09 21:03 A. Camsonne Electron detector seems still alive !
  • 287562 09/02/09 20:51 xiaohui Left Q3 strip
  • 287561 09/02/09 20:37 Zafar / Bob reducing gain DID help
  • 287560 09/02/09 20:26 Fatiha Q3 left arm tripped
  • 287557 09/02/09 20:17 R. Michaels Pre-Gain did not change Cavities
  • 287555 09/02/09 20:08 R. Michaels Cavity Gain adjustments
  • 287554 09/02/09 20:04 Fatiha Beam back
  • 287552 09/02/09 20:00 Zafar/Bob cavities vs striplines for stable beam.
  • 287551 09/02/09 19:56 Fatiha Beam down again
  • 287548 09/02/09 19:38 A. Camsonne Compton Analog board on upper half of B and C unplugged
  • 287547 09/02/09 19:33 Fatiha Asked for beam back
  • 287546 09/02/09 19:24 Zafar/Bob Cavities and Striplines in RHRS
  • 287545 09/02/09 19:04 xiaohui BPM4AX back to normal
  • 287544 09/02/09 19:04 Fatiha Hall going to controlled access
  • 287541 09/02/09 18:35 Aniol target JT valve adjustement
  • 287539 09/02/09 18:10 Apar BCM 4AX off 0
  • 287538 09/02/09 18:02 Fatiha Beam down since a while
  • 287536 09/02/09 17:43 dalton Stuck bit in CMU DAQ scaler
  • 287535 09/02/09 17:18 jpchen reduce parameter dE/dx further down
  • 287534 09/02/09 17:08 silwal boiling tests, Tuesday 09/01/09, swing
  • 287533 09/02/09 17:03 jpchen Target: coolant for high current (100 uA) running
  • 287530 09/02/09 16:55 jpchen Changed HP heater parameter to reduce false alarm
  • 287529 09/02/09 16:55 Fatiha Feedback Monitor
  • 287528 09/02/09 16:53 Fatiha Beam current screen
  • 287527 09/02/09 16:52 R. Michaels automated ABU turned on
  • 287526 09/02/09 16:51 Fatiha Hall A General tools
  • 287525 09/02/09 16:20 Higinbotham Shift Summary
  • 287524 09/02/09 16:19 Higinbotham Hanging Raster Cable
  • 287523 09/02/09 16:19 O.Glamazdin Moller spin dance details
  • 287521 09/02/09 16:03 W. Deconinck Target screen shots
  • 287518 09/02/09 15:56 Jeong Han Hall A General Tools
  • 287517 09/02/09 15:03 mercado Boiling Studies - Current Scan - 40 to 100uA
  • 287512 09/02/09 14:17 Z. Ahmed. Integration gate time changed.
  • 287510 09/02/09 14:02 R. Michaels BPM cavities (part IV).
  • 287508 09/02/09 13:55 R. Michaels Cavity BPMs (Part III of check)
  • 287505 09/02/09 13:37 R. Michaels cavities vs bcm1
  • 287502 09/02/09 13:32 mercado ATTENTION Shift Workers!
  • 287501 09/02/09 13:30 R. Michaels Cavity checkout at 100uA (part I)
  • 287496 09/02/09 13:13 R. Michaels HapNoPvdis crate now
  • 287492 09/02/09 12:59 Higinbotham Energies Around The Machine
  • 287491 09/02/09 12:56 Jeong Han Stop run 11879
  • 287489 09/02/09 12:37 dalton New lumi high voltages for 100 uA running
  • 287486 09/02/09 12:16 W. Deconinck 100 uA is pushing up cryo inlet temp to 15.1 K
  • 287484 09/02/09 12:09 M Friend Compton Spectrum
  • 287483 09/02/09 12:04 dalton VNC HV connection hangs up twice
  • 287480 09/02/09 11:39 Jeong Han Promt anlaysis
  • 287477 09/02/09 11:09 W. Deconinck Target heater and cryo inlet temperature strip charts
  • 287476 09/02/09 11:08 W. Deconinck Target cryo settings at 100 uA
  • 287474 09/02/09 11:07 Higinbotham 100 uA 25 cm LH2
  • 287473 09/02/09 10:54 W. Deconinck Trying to get higher cryo flow
  • 287472 09/02/09 10:03 W. Deconinck Maximum e current
  • 287471 09/02/09 08:54 dalton Compton end of run failed
  • 287465 09/02/09 08:02 Kalyan Shift Summary
  • 287463 09/02/09 07:53 hoyoung Hall A general tools
  • 287462 09/02/09 07:37 silwal made controlled access in the right arm, found that led was turned off, relay switch unplugged
  • 287461 09/02/09 07:37 R. Michaels defaults for GM
  • 287458 09/02/09 07:08 silwal rebooting creates, ADC pedestel set
  • 287449 09/02/09 05:40 A. Camsonne Removed transversity directory on /adaql10/work1
  • 287446 09/02/09 05:26 Kalyan Hall A tools
  • 287445 09/02/09 05:15 silwal space in /adaql10/work1/
  • 287444 09/02/09 04:47 A. Camsonne Terabyte array
  • 287441 09/02/09 04:14 silwal ROOTfiles -> /adaql5/work1/happex3/ROOTfiles full
  • 287438 09/02/09 02:53 Kalyan Target temperature and pressure strip charts
  • 287437 09/02/09 02:52 Kalyan Target screen
  • 287436 09/02/09 02:44 hoyoung Run #11864 is junk
  • 287431 09/02/09 01:33 Kalyan Raster size 5x4.5 for the production
  • 287430 09/02/09 01:33 silwal det1 (LHRS) with base H3/Bugs Bunny vs det2 (RHRS) HC/Daffy Duck signal, counting DAQ
  • 287425 09/02/09 01:02 hoyoung Lumi voltage change
  • 287422 09/02/09 00:50 silwal changed lumi HV for 50 uA beam
  • 287419 09/02/09 00:33 hoyoung Magnet strips
  • 287418 09/02/09 00:32 hoyoung Ion chamber strip
  • 287417 09/02/09 00:32 hoyoung Beam position and current strip
  • 287416 09/02/09 00:31 hoyoung Hall A general tools
  • 287411 09/02/09 00:00 Kai Pan swing shift summay
  • Tuesday

  • 287410 09/01/09 23:49 Fatiha Run list
  • 287409 09/01/09 23:49 Kai reason for no data taking from 22:45 to 23:30
  • 287408 09/01/09 23:45 aniol target power and crystat
  • 287407 09/01/09 23:44 aniol target screen
  • 287405 09/01/09 23:33 Kai snapshot of Hall A general tool
  • 287404 09/01/09 23:30 Kai keep taking data
  • 287393 09/01/09 21:34 A. Camsonne Compton Electron detector status
  • 287391 09/01/09 21:14 A. Camsonne Compton Electron detector noise
  • 287390 09/01/09 21:14 Kai raster scan
  • 287389 09/01/09 21:06 Fatiha Spot++ 11.0x8.5 (MCC Units), 5x4.5 our units
  • 287387 09/01/09 21:00 Fatiha Spot++ 7.7 x 8 (MCC units), 3.5 x 4.5 at BPM b ( our units)
  • 287384 09/01/09 20:54 Fatiha Spot ++: beam current = 5 uA raster size 7.7 x 6 ( MCC), 3.5 x 3.5 (ours)
  • 287380 09/01/09 20:38 silwal PITA slope is -4081, charge feedback back on
  • 287377 09/01/09 20:18 silwal lumi voltages changed for 100 uA beam
  • 287374 09/01/09 19:56 A. Camsonne Compton electron detector inhibit logic
  • 287371 09/01/09 18:57 Kai MCC cannot reach 80uA right now
  • 287370 09/01/09 18:52 Kai Detector (VDC) Flammable Gas Sensor is alarming
  • 287369 09/01/09 18:51 Fatiha Ion Chamber.stp
  • 287368 09/01/09 18:48 Fatiha More snapshots
  • 287367 09/01/09 18:46 Fatiha Hall A Genaral tools
  • 287365 09/01/09 18:11 Fatiha Spot++
  • 287361 09/01/09 18:05 Kai beam back
  • 287360 09/01/09 17:46 M Friend Compton FADC Clock
  • 287358 09/01/09 17:20 R. Michaels L-HRS Qsq
  • 287353 09/01/09 16:05 Jeong Han screenshots for GUIs
  • 287352 09/01/09 16:05 Jeong Han screenshots for GUIs
  • 287350 09/01/09 16:04 Ole Hansen Day Shift Summary
  • 287349 09/01/09 15:26 Ole Hansen Beam taken away
  • 287348 09/01/09 15:16 Ole Hansen 10uA on lH2
  • 287347 09/01/09 15:01 Ole Hansen Spot ok
  • 287346 09/01/09 15:00 W. Deconinck Raster from spot++
  • 287341 09/01/09 14:45 Ole Hansen S0 right off
  • 287340 09/01/09 14:41 Ole Hansen S0 detectors on
  • 287339 09/01/09 14:37 Ole Hansen Beam restore
  • 287338 09/01/09 14:32 Ole Hansen Access finished, preparing for beam
  • 287337 09/01/09 12:35 M Friend Correction
  • 287336 09/01/09 11:18 M Friend Compton: Lead Filter Replaced
  • 287335 09/01/09 09:51 Ole Hansen Ramping down magnets
  • 287334 09/01/09 08:59 Ole Hansen PSS problem
  • 287331 09/01/09 08:40 R. Michaels stage disk server is back
  • 287330 09/01/09 08:33 Jeong Han Screenshots for guis
  • 287329 09/01/09 08:27 silwal owl shift summary
  • 287328 09/01/09 08:12 Jeong Han Run with charge feedback off
  • 287326 09/01/09 08:06 R. Michaels staging disk is down
  • 287325 09/01/09 08:04 Abdurahim target main screen
  • 287323 09/01/09 08:04 R. Michaels pan DB mistake
  • 287322 09/01/09 08:03 Abdurahim The target strip charts at the end of Owl
  • 287321 09/01/09 08:02 Abdurahim Target strip charts at the of the Owl
  • 287320 09/01/09 07:49 hoyoung Hall A general tools screen
  • 287319 09/01/09 07:31 R. Michaels re: ADCX6 err messages (false ?)
  • 287313 09/01/09 06:10 silwal pan analysis errors
  • 287312 09/01/09 06:10 silwal pan analysis errors
  • 287311 09/01/09 05:59 Abdurahim Target main screen
  • 287310 09/01/09 05:58 Abdurahim Target strip charts
  • 287307 09/01/09 04:55 hoyoung Charge feedback enabled
  • 287304 09/01/09 04:02 Abdurahim target main screen snap shot
  • 287303 09/01/09 04:00 Abdurahim target strip charts
  • 287302 09/01/09 03:59 silwal/hoyou BCM asymmetries close to zero
  • 287301 09/01/09 03:49 silwal/hoyou charge feedback turned off
  • 287296 09/01/09 02:34 silwal HV for detectors, new HV for LHRS
  • 287295 09/01/09 02:20 silwal relay switch back in LHRS
  • 287294 09/01/09 01:59 Abdurahim Target strip charts
  • 287292 09/01/09 01:56 Abdurahim Target Motion
  • 287291 09/01/09 01:38 mcnulty Alignment Plots from Latest H2 running
  • 287290 09/01/09 01:14 mcnulty Happex Detector Alignment Plots
  • 287288 09/01/09 00:53 Abdurahim Target Motion
  • 287286 09/01/09 00:10 hoyoung Magnet strips
  • 287285 09/01/09 00:08 hoyoung Ion chambers strip
  • 287284 09/01/09 00:08 hoyoung Beam position and current
  • 287283 09/01/09 00:07 hoyoung hall A general tools
  • 287281 09/01/09 00:01 target end of swing shift, all good!
  • 287280 09/01/09 00:00 target end of swing shift, all good!